### v2

# Resurgent structure of 2d Yang-Mills theory on a torus

Abstract
We study the resurgent structure of the topological string dual to 2d $U(N)$ Yang-Mills on torus. We find closed form formulas for instanton amplitudes up to arbitrarily high instanton orders, based on which we propose the non-perturbative partition function including contributions from all the real instantons, which is real for positive modulus and string coupling. We also explore complex instantons and find two infinite towers of them. We expect them to correspond to BPS states in type II string.
Keywords:
2d Yang-Mills, topological string, resurgence, Stokes transformation, non-perturbative, BPS states
preprint: USTC-ICTS/PCFT-25-29
1 Introduction
Two dimensional $U(N)$ Yang-Mills theory is an interesting non-supersymmetric QFT which is exactly solvable yet still non-trivial Migdal:1975zg ; Rusakov:1990rs ; Witten:1991we ; Witten:1992xu . [^1] See Blau:1993hj ; Cordes:1994fc for instance for good reviews. Furthermore, it is known that in the large $N$ limit, it is dual to a string theory Gross:1992tu ; Gross:1993hu ; Gross:1993yt . The partition function of the 2d Yang-Mills factorizes to the product of a chiral part and an anti-chiral part, and the chiral part can be interpreted as counting holomorphic maps from the string worldsheet to the target space of the gauge theory. However, the true nature of the string theory is still mysterious when the target space is a generic Riemann surface of arbitrary genus and when the ’t Hooft coupling is finite. See Cordes:1994sd ; Horava:1995ic ; Vafa:2004qa ; Aganagic:2004js ; Aharony:2023tam ; Benizri:2025xmz ; Komatsu:2025sqo for recent development.
When the target space is a torus, the situation is much more clear. By relating partition function of the 2d Yang-Mills with that of 4d BPS black holes and invoking the OSV conjecture Ooguri:2004zv , the dual string theory is identified as topological string theory put on a double line bundle over an elliptic curve Vafa:2004qa . For instance, the partition function of the 2d Yang-Mills theory factorizes in the large $N$ limit, and the chiral component is argued to be the partition function of the dual topological string theory. Indeed, it was shown in Dijkgraaf:1996iy ; Okuyama:2019rqn ; Huang:2021zko that the chiral partition function of 2d Yang-Mills can be promoted to non-holomorphic functions which satisfy the BCOV holomorphic anomaly equations Bershadsky:1993ta ; Bershadsky:1993cx , a hallmark of the partition function of topological string theory.
However, it is not known how to formulate the large $N$ duality precisely when $N$ is finite. At large but finite $N$, the factorization is only schematic and it needs to be modified by summing over RR fluxes over the finite $T^{2}$ Vafa:2004qa ; Dijkgraaf:2005bp . More importantly, while on the one side the partition function of 2d Yang-Mills is non-perturbatively well-defined for any finite $N$, on the other side the topological string partition function is only defined perturbatively. One therefore has to find its non-perturbative completion, taking into account appropriate non-perturbative corrections of the order $\mathcal{O}({\rm e}^{-N})$. There have been various attempts to find these non-perturbative corrections Lelli:2002gr ; Matsuo:2004nn ; deMelloKoch:2005rq ; Dhar:2006ru ; Okuyama:2018clk . The most promising proposal is given by Okuyama and Sakai Okuyama:2018clk , based on free fermion formulation of the 2d Yang-Mills Minahan:1993np ; Douglas:1993wy . However, only the 1-instanton amplitude of the proposal has been tested, and its multi-instanton amplitudes are not only un-tested, but in fact suffer from some limitations.
In this paper, we make a new proposal for the full non-perturbative partition function for the topological string dual to chiral 2d Yang-Mills, which overcomes the limitations that plague previous proposals. Our proposal exploits recent studies of non-perturbative corrections to topological string partition functions by the powerful method of resurgence theory Ecalle .  [^2] Resurgence theory has been used to study 2d Yang-Mills for small $N$ Fujimori:2022qij . See Griguolo:2024ecw for perturbative expansion at instanton sectors for 2d Yang-Mills with small $N$.
The resurgence theory is a powerful mathematical framework that uncovers the intimate relationship between asymptotic perturbative series and its non-perturbative corrections, which is crucial for making sense of the asymptotic perturbative series itself. See Sauzin:2007resurgent ; Marino:2012zq ; Aniceto:2018bis for useful reviews. The resurgence theory claims that in order to incorporate appropriately non-perturbative corrections, one needs to upgrade the perturbative power series to trans-series which are subject to the same set of constraints or equations. Furthermore, at least a subset of the non-perturbative corrections are closely related to the perturbative series, and they transform to each other via Stokes transformations. These non-perturbative corrections are said to be in the resurgent structure of the perturbative series Gu:2021ize . They are the most important non-perturbative corrections, and are usually indispensable if one wishes to make sense of the asymptotic perturbative series.  [^3] There could be non-perturbative corrections in different resurgent structures than that of the perturbative series, as in the double-well and the cosine model of QM, explained for instance in Dunne:2014uwm .
In the case of topological string theory, the resurgence theory has been applied heavily to uncover non-perturbative corrections to the partition function. In particular, it has been proposed Couso-Santamaria2016 ; Couso-Santamaria2015 ; Couso-Santamaria:2016vwq that non-perturbative partition function of topological string, in the form of a trans-series, also satisfies the BCOV holomorphic anomaly equations which control the perturbative partition function. In addition, when solving the non-perturbative partition function from the holomorphic anomaly equations, the integration constant, known as the holomorphic ambiguity, can be fixed completely by exploiting the Stokes transformation that connect the perturbative partition function and its non-perturbative corrections. Here we restrict ourselves to the most important non-perturbative corrections, those in the resurgent structure of the perturbative partition function. Important progress has been made in this direction for topological string theory with local Calabi-Yau threefolds Gu:2022sqc or simple compact Calabi-Yau threefolds Gu:2023mgf as background geometry.  [^4] See further development in Marino:2023gxy ; Iwaki:2023rst ; Marino:2023nem ; Alexandrov:2023wdj ; Gu:2024wag and Marino:2024tbx for a review. Full non-perturbative corrections for both single and multi-instanton amplitudes are solved up to arbitrary instanton orders, and the Stokes transformation of the perturbative partition function in terms of these instanton amplitudes are studied. An important feature of these instanton amplitudes is that they are functionals of the perturbative partition function whose moduli are shifted discretely, which is the typical effect of D-brane insertion Aganagic:2003qj ; Aganagic:2011mi . In fact, these instanton amplitudes are conjectured to be closely related to BPS states of stable bound states of D-branes in type IIA string on the same background geometry. The action of the instanton amplitude is the central charge of a BPS state Drukker:2011zy ; Couso-Santamaria2016 ; Couso-Santamaria2015 , and the integer Stokes constant that characterizes the associated Stokes transformation is the BPS multiplicity, first suggested in Gu:2021ize ; Gu:2022sqc ; Gu:2023mgf with supporting evidence appearing later in Iwaki:2023rst ; Gu:2023wum ; Douaud:2024khu ; Marino:2024yme . See Bridgeland:2016nqw ; Bridgeland:2017vbr ; Alim:2021mhp ; Bridgeland:2024ecw ; Alim:2024dyi ; Alim:2015qma ; Couso-Santamaria2017a ; Li:2025zyr also for related development.
In this paper, we apply the same method to find non-perturbative corrections to the partition function of the topological string dual to 2d Yang-Mills. We find closed form formulas for both single and multi-instanton amplitudes up to arbitrary instanton orders, and they turn out to be slightly different from the formulas in the usual topological string Gu:2022sqc ; Gu:2023mgf . Our 1-instanton amplitude agrees with that of Okuyama and Sakai Okuyama:2018clk , but our multi-instanton amplitudes are different. They pass high precision numerical tests from resurgence theory, and they overcome the limitations suffered by previous proposals. In particular, our formulas are valid for non-vanishing $\theta$ as well. Based on these results, we propose the non-perturbative partition function,
$\mathcal{Z}^{\text{np}}(t;g_{s};\bm{\sigma})=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}\pm\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,2,\ldots,n}\right)\mathscr{S}^{(\pm)}\mathcal{Z}(t+ng_{s};g_{s}),$ (1.1)
taking into account contributions from all the real instantons, and it is real when both the modulus $t$ and the string coupling $g_{s}$ are positive. It involves an infinite family of real parameters $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots)$, and in the simplest case $\bm{\sigma}=\bm{0}$, it is just the medium resummation of the perturbative partition function. Here $B_{n}(\ldots)$ are complete Bell polynomials, and $\mathscr{S}^{(\pm)}$ denote the lateral Borel resummations. Note that one can choose either of the two lateral Borel resummations $\mathscr{S}^{(\pm)}$ (see (4.10) for definition) with the same sign $\pm$ in the Bell polynomials to match it. We believe this should be used to formulate the precise duality between 2d Yang-Mills on torus and topological string for finite $N$.
In addition, we also explore complex instantons, and find two infinite towers of complex instanton sectors. They should correspond to non-trivial BPS D-brane bound states in type II string. We also discuss a little the wall-crossing behavior of these instanton sectors.
The structure of the paper is as follows. In Section 2 we review the string interpretation of 2d $U(N)$ Yang-Mills on torus in the large $N$ limit, the perturbative partition function of the dual topological string, and the proposal of Okuyama and Sakai Okuyama:2018clk to make the latter non-perturbative. We comment on the limitations of their proposal. In Section 3 we make the duality between the 2d $U(N)$ Yang-Mills on torus and topological string more concrete by spelling out the special geometry relations and the singular points in the moduli space, which are crucial for formulating precisely the holomorphic anomaly equations that control the dual topological string theory. Section 4 contains the main results of this paper. We work out the instanton amplitudes, derive the real non-perturbative partition function, and finally explore complex instanton sectors. We conclude in Section 5.
2 Perturbative and non-perturbative 2d Yang-Mills
2.1 String interpretation
Consider two dimensional Yang-Mills theory with gauge group $G=U(N)$ on a torus $T^{2}$ with $\theta$ angle turned on. The action of the theory is
$S=\frac{1}{2g^{2}_{\text{YM}}}\int_{T^{2}}\operatorname{tr}(F\wedge\star F)+\theta\operatorname{tr}F.$ (2.1)
Here $\operatorname{tr}$ is the trace in the fundamental representation. The partition function can be calculated by summing over representations $R$ of the gauge group Migdal:1975zg ; Rusakov:1990rs
$Z_{N}^{\text{YM}}=\sum_{R}\exp\left(-\frac{1}{2}g_{\text{YM}}^{2}C_{2}(R)+{\mathsf{i}}\,\theta\,C_{1}(R)\right),$ (2.2)
where $C_{1}(R),C_{2}(R)$ denote the first and the second Casimirs of the representation.
It was argued in Gross:1992tu ; Gross:1993hu that in the large $N$ limit,
$N\rightarrow\infty,\quad g_{\text{YM}}\rightarrow 0,\quad Ng_{\text{YM}}^{2}\;\text{fixed},$ (2.3)
the partition function $Z_{N}^{\text{YM}}$ factorizes to a product of chiral and anti-chiral components, and the chiral component can be interpreted as the partition function of a string theory. By relating to 4d BPS black hole partition function and applying the OSV conjecture Ooguri:2004zv , it was pointed out in Vafa:2004qa that the string theory is none other than the topological string theory.  [^5] This is only true when the 2d Yang-Mills is put on a torus. When the target space is a generic Riemann surface of arbitrary genus and the ’t Hooft coupling is finite, the exact string theory that corresponds to the chiral partition function is not known. See for instance Cordes:1994fc ; Horava:1995ic ; Aganagic:2004js ; Komatsu:2025sqo ; Benizri:2025xmz for works in this direction. More explicitly, the gauge theory partition function factorizes by
$Z_{N}^{\text{YM}}=|\mathcal{Z}_{\text{top}}(t;g_{s})|^{2}.$ (2.4)
where $\mathcal{Z}_{\text{top}}(t;g_{s})$ is the partition function of the topological string theory whose target space is the total space of the double line bundle over the torus $E$
$X_{E}:\;\mathcal{O}(1)\oplus\mathcal{O}(-1)\rightarrow E,$ (2.5)
and which is described by the Gromov-Witten theory of an elliptic curve dijkgraaf:1995mirror . The string coupling $g_{s}$ and the complexified Kähler modulus $t$ are given respectively by
$g_{s}=g_{\text{YM}}^{2}A,\quad t=\frac{1}{2}Ng_{s}-{\mathsf{i}}\theta,$ (2.6)
where $A$ is the torus area, which we will set to one. The free energy then has the genus expansion
$\mathcal{F}(t;g_{s})=\log\mathcal{Z}_{\text{top}}(t;g_{s})=\sum_{g\geq 0}g_{s}^{2g-2}\mathcal{F}_{g}(t).$ (2.7)
The first few terms are
$\mathcal{F}_{0}(t)=-\frac{t^{3}}{6},\quad\mathcal{F}_{1}(t)=-\log\eta(q)=\frac{t}{24}-\log\prod_{n=1}^{\infty}(1-q^{n}),$ (2.8)
with
$q={\rm e}^{-t}={\rm e}^{2\pi{\mathsf{i}}\tau},\quad\tau={\mathsf{i}}t/(2\pi).$ (2.9)
For most part of the paper, we will focus on the region with $\operatorname{Re}t>0$, i.e. $\operatorname{Im}\tau>0$.
2.2 Perturbative free energies
The free energy has very nice properties. First of all, it was shown in dijkgraaf:1995mirror and proved in Kaneko1995:msc that the free energy $\mathcal{F}_{g}(t)$ with $g\geq 2$ is a quasi-modular form of weight 6g66g-6 given by a combination of Eisenstein series. Second, the free energy $\mathcal{F}_{g}(t)$ can be calculated recursively Okuyama:2018clk . Let us split the holomorphic perturbative partition function via
$\mathcal{Z}^{\text{top}}(t,g_{s})=\mathcal{Z}_{01}(t,g_{s})\widehat{\mathcal{Z}}^{\text{top}}(t,g_{s}),$ (2.10)
where
$\mathcal{Z}_{01}(t,g_{s})=\exp\left(\mathcal{F}_{0}/g_{s}^{2}+\mathcal{F}_{1}\right),$ (2.11)
and
$\widehat{\mathcal{Z}}^{\text{top}}(t,g_{s})=\exp\Bigg(\sum_{g\geq 2}g_{s}^{2g-2}\mathcal{F}_{g}(t)\Bigg)=\sum_{n=0}^{\infty}g_{s}^{2n}\mathcal{Z}_{n}^{\text{top}}(t),$ (2.12)
with $\mathcal{Z}_{0}^{\text{top}}(t)=1$. Then $\mathcal{Z}_{n}^{\text{top}}(t)$ satisfies the recursive relation Okuyama:2018clk
$\mathcal{Z}_{n}^{\text{top}}(t)=h_{n}^{\text{top}}-\sum_{m=1}^{n}\frac{[D_{-1}+\frac{1}{3}D_{3}]^{2m}}{(2m)!}\mathcal{Z}_{n-m}^{\text{top}}(t)\cdot(2D_{3})^{m}1.$ (2.13)
Here $h_{n}^{\text{top}}$ is defined by the generating series
$\sum_{n=0}^{\infty}g_{s}^{2n}h_{n}^{\text{top}}=\exp\left(\sum_{\ell=1}^{\infty}g_{s}^{2\ell}\frac{e_{\ell}}{(2\ell)!}\right),$ (2.14)
where
$e_{\ell}=\frac{B_{2\ell+2}}{2\ell+2}e^{-2\ell}D^{2\ell-1}E_{2\ell+2}(q),$ (2.15)
and
$D=q\frac{\partial}{\partial q}.$ (2.16)
In addition the derivatives $D_{k}$ are defined by
$D_{k}=D+\frac{kE_{2}}{24},$ (2.17)
and we impose in the recursion relation (2.13) that $D_{-1}$ only acts on $\mathcal{Z}^{\text{top}}_{*}$ while $D_{3}$ only acts on $1$. As an example, we find
$\mathcal{Z}^{\text{top}}_{1}(t)=\frac{5E_{2}^{3}-3E_{2}E_{4}-2E_{6}}{51840}.$ (2.18)
The free energy $\mathcal{F}_{g}(t)$ can then be calculated from the exponential relation (2.12).
To compute the free energy $\mathcal{F}_{g}(t)$ efficiently, we begin with a high-order, e.g., 2000th order, $q$-series expansion of $e_{\ell}$. Utilizing the recursive relation
$h_{n}^{\text{top}}=\frac{1}{n}\sum_{l=1}^{n}h_{n-l}^{\text{top}}\frac{l\,e_{l}}{(2l)!},$ (2.19)
derived from (2.14), the $q$-series expression of $h_{n}^{\text{top}}$ can be computed efficiently. Substituting $h_{n}^{\text{top}}$ into the recursive relation (2.13), we are able to obtain the 2000th-order $q$-series expression of $\mathcal{Z}_{n}^{\text{top}}(t)$ for $n\leq 200$. Since $\mathcal{Z}_{n}^{\text{top}}(t)$ is finitely generated by Eisenstein series, a sufficiently high-order $q$-series expansion allows one, in principle, to recover the modular expression of $\mathcal{Z}_{n}^{\text{top}}(t)$.
Finally, the free energy $\mathcal{F}_{g}(t)$ with $g\geq 1$ of topological string theory has the property that it can be uplifted to a non-holomorphic version $F_{g}(t,\bar{t})$, which satisfies the BCOV holomorphic anomaly equations Bershadsky:1993cx ; Bershadsky:1993ta . In the case of chiral 2d Yang-Mills, the uplift can be achieved by considering the bosonization of the free fermion formulation Douglas:1993wy ; Dijkgraaf:1996iy . It was emphasized in Okuyama:2019rqn that in this process, one should distinguish two kinds of $E_{2}$, those coming from period integrals over the torus and those from the so-called propagator, and the uplift is done via replacing the $E_{2}$ of the latter type by the almost holomorphic Eisenstein series
$\widehat{E}_{2}(\tau,\bar{\tau})=E_{2}(\tau)-\frac{3}{\pi\tau_{2}},$ (2.20)
where $\tau_{2}=\operatorname{Im}\tau$ is the imaginary part of $\tau$, but not the $E_{2}$ of the former type. This is different from the usual case in topological string theory where all $E_{2}$ are replaced by $\widehat{E}_{2}$ when uplifting holomorphic free energy to non-holomorphic free energy Huang:2006si ; Huang:2013yta .  [^6] In the chiral 2d Yang-Mills theory, if all $E_{2}$ are replaced by $\widehat{E}_{2}$, the resulting non-holomorphic free energies do not satisfy holomorphic anomaly equations Okuyama:2019rqn . Once this is taken care of, it was checked in Okuyama:2019rqn and later proved in Huang:2021zko that indeed the non-holomorphic free energies satisfy the holomorphic anomaly equations
$\partial_{S}F_{g}=\frac{1}{2}\left(\mathcal{D}^{2}F_{g-1}+\sum_{h=1}^{g-1}\mathcal{D}F_{h}\mathcal{D}F_{g-h}\right),\quad g\geq 2.$ (2.21)
Here on the right hand side, $\mathcal{D}$ is the covariant derivative on the moduli space acting on the free energies via
$\mathcal{D}F_{g}=DF_{g},\;\;\mathcal{D}^{2}F_{g}=(D+S)DF_{g}.$ (2.22)
On the left hand side, $\partial_{S}$ is the derivative with respect to the propagator defined by
$S=\frac{1}{t+\bar{t}}=\frac{1}{4\pi\tau_{2}}=\frac{E_{2}-\widehat{E}_{2}}{12}.$ (2.23)
The derivation is equivalent to the derivative with respect to $\bar{\tau}$ via
$\partial_{S}=8\pi{\mathsf{i}}(\tau_{2})^{2}\partial_{\bar{\tau}}.$ (2.24)
The genus one free energy is not subject to the holomorphic anomaly equations (2.21) and its non-holomorphic uplift is
$F_{1}=\log\frac{\sqrt{2\pi S}}{\eta}.$ (2.25)
In fact, it serves as the initial condition for solving the holomorphic anomaly equations. For instance, by direct integration, we find
$F_{2}=\frac{5}{24}S^{3}-\frac{E_{2}}{48}S^{2}+\frac{2E_{4}-E_{2}^{2}}{1152}S+\mathcal{F}_{2}(t).$ (2.26)
The holomorphic part $\mathcal{F}_{2}(t)$ is the integration constant, also known as the holomorphic ambiguity, and it can be obtained from (2.12) and (2.18),
$\mathcal{F}_{2}(t)=\mathcal{Z}_{1}^{\text{top}}(t)=\frac{5E_{2}^{3}-3E_{2}E_{4}-2E_{6}}{51840}.$ (2.27)
In general, the non-holomorphic free energy $F_{g}(t,\bar{t})$ is a polynomial in $S$ of degree 3g33g-3
$F_{g}=\sum_{k=0}^{3g-3}P_{n}(t)S^{n},$ (2.28)
where the holomorphic part $P_{0}(z)$ is the holomorphic free energy $\mathcal{F}_{g}(t)$.
2.3 Non-perturbative proposal
It was argued in Vafa:2004qa that the factorization (2.4) is only schematic and to make it exact one has to sum over RR-fluxes over the torus, arriving at
$Z_{N}^{\text{YM}}=\sum_{n\in\mathbb{Z}}\mathcal{Z}_{\text{top}}(t+ng_{s};g_{s})\overline{\mathcal{Z}}_{\text{top}}(t-ng_{s};g_{s}).$ (2.29)
An even more refined version is proposed in Dijkgraaf:2005bp . It is, however, difficult to generalize formulas like (2.29) to cases with large but finite $N$. As argued in Okuyama:2018clk , in this case the meaning of (2.29) is unclear: the left hand side is well-defined for finite $N$ and thus finite $g_{s}$, while the right hand side is only defined as asymptotic perturbative power series of $g_{s}$.
An attempt to solve this problem was made in Okuyama:2019rqn , in which process, a non-perturbative version of topological string free energy was proposed. The partition function of the 2d Yang-Mills theory allows a free fermion formulation Minahan:1993np ; Douglas:1993wy
$Z_{N}^{\text{YM}}=$ $\oint\frac{{\rm d}x}{2\pi{\mathsf{i}}x^{N+1}}\prod_{n\in\mathbb{Z}+\frac{N-1}{2}}\left(1+x{\rm e}^{{\mathsf{i}}n\theta}p^{\frac{1}{2}n^{2}}\right)$
$=$ $\oint\frac{{\rm d}x}{2\pi{\mathsf{i}}x^{N+1}}\exp\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}x^{k}\begin{cases}\vartheta_{2}({\rm e}^{{\mathsf{i}}k\theta},p^{k}),\quad&\text{even }N,\\ \vartheta_{3}({\rm e}^{{\mathsf{i}}k\theta},p^{k}),\quad&\text{odd }N,\end{cases}$ (2.30)
where $p={\rm e}^{-g_{s}}$. Let us focus on the simple scenario $\theta=0$ and assume that $N$ is always even. The partition function reduces to
$Z_{N}^{\text{YM}}=\oint\frac{{\rm d}x}{2\pi{\mathsf{i}}x^{N+1}}\prod_{n\in\mathbb{Z}+\frac{1}{2}}\left(1+xp^{\frac{1}{2}n^{2}}\right).$ (2.31)
It can be written as
$Z_{N}^{\text{YM}}=\sum_{\begin{subarray}{c}N_{+}+N_{-}=N\\ N_{\pm}\geq 0\end{subarray}}\mathcal{Z}_{N_{+}}\overline{\mathcal{Z}}_{N_{-}},$ (2.32)
where $\mathcal{Z}_{N_{+}}$ and $\mathcal{Z}_{N_{-}}$ are defined via the generating series
$\sum_{N_{+}=0}^{\infty}\mathcal{Z}_{N_{+}}x^{N_{+}}=\prod_{n\in\mathbb{Z}_{\geq 0}+1/2}\left(1+xp^{n^{2}/2}\right),$ (2.33a)
$\sum_{N_{-}=0}^{\infty}\overline{\mathcal{Z}}_{N_{-}}x^{N_{-}}=\prod_{n\in-\mathbb{Z}_{\geq 0}-1/2}\left(1+xp^{n^{2}/2}\right).$ (2.33b)
Eq. (2.32) looks similar to (2.29), and it was proposed in Okuyama:2018clk that $\mathcal{Z}_{N_{+}}$ be regarded as the non-perturbative completion of topological string partition function,
$\mathcal{Z}^{\text{OS}}(t,g_{s})=\mathcal{Z}_{N_{+}},$ (2.34)
with the dictionary $t=N_{+}g_{s}$, in the sense that
$\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})\left(1+\mathcal{O}({\rm e}^{-1/g_{s}})\right).$ (2.35)
More explicitly, one can show that Okuyama:2018clk
$\mathcal{Z}^{\text{OS}}(t,g_{s})=\sum_{n=0}^{\infty}\phi_{n}(p)\mathscr{S}^{(+)}\mathcal{Z}^{\text{top}}(t+ng_{s};g_{s}),$ (2.36)
where $\mathscr{S}^{(+)}$ is the Borel resummation (see Sec. 4.1) and the functions $\phi_{k}$ are defined via the generating series
$\sum_{n=0}^{\infty}\phi_{n}(p)y^{n}=\exp\left(\frac{{\mathsf{i}}}{2}\sum_{k=0}^{\infty}\frac{(-y)^{k}}{k}\vartheta_{2}(p^{k})\right).$ (2.37)
In the leading instanton order
$\phi_{1}(p)=-\frac{{\mathsf{i}}}{2}\vartheta_{2}(p),$ (2.38)
and we have
$\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})-\frac{{\mathsf{i}}}{2}\vartheta_{2}(q)\mathcal{Z}^{\text{top}}(t+g_{s};g_{s})+\ldots$ (2.39)
Since in the $g_{s}\rightarrow 0$ limit
$\vartheta_{2}(p)=\sqrt{\frac{2\pi}{g_{s}}}\vartheta_{4}({\rm e}^{-4\pi^{2}/g_{s}})\approx\sqrt{\frac{2\pi}{g_{s}}},$ (2.40)
we further have the approximation,
$\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})-\frac{{\mathsf{i}}}{2}\sqrt{\frac{2\pi}{g_{s}}}\mathcal{Z}^{\text{top}}(t+g_{s};g_{s})+\ldots.$ (2.41)
Here $\mathcal{Z}^{\text{top}}(t+g_{s};g_{s})/\mathcal{Z}^{\text{top}}(t;g_{s})\sim{\rm e}^{-t^{2}/2g_{s}}$ is of non-perturbative order, representing corrections from 1-instanton amplitude. Eq. (2.41) has been verified in detail by resurgence analysis Okuyama:2018clk .
However, there are several problems with this proposal. First of all, even though this proposal includes corrections from multi-instanton amplitudes to arbitrary high orders, only the 1-instanton amplitude was tested Okuyama:2018clk . Second, this proposal was derived with the assumption that $N$ is even and $\theta=0$, which is simply too restrictive. On the one hand, when $N$ is odd, (2.30) seems to indicate that in the generating series (2.37) $\vartheta_{2}$ should be replaced by $\vartheta_{3}(p)$, which nonetheless cannot be distinguished by the 1-instanton test, as in the $g_{s}\rightarrow 0$ limit,
$\vartheta_{2}(p)\approx\vartheta_{3}(p)\approx\sqrt{\frac{2\pi}{g_{s}}}.$ (2.42)
On the other hand, it was admitted in Okuyama:2018clk that their proposal runs into problems for non-vanishing $\theta$. Finally, as we will see in Section 4, the proposed non-perturbative partition function has a fictitious non-vanishing imaginary part when $t$ and $g_{s}$ are positive. We will overcome these problems and make a more consistent proposal for the non-perturbative partition function.
3 Special geometry of chiral 2d Yang-Mills
In this section, we discuss some important properties of the moduli space of topological string theory (see klemm2018b for instance for review) and verify that the chiral 2d Yang-Mills theory indeed enjoys these properties. We restrict our discussion to the cases where the moduli space is complex one dimensional, which is parametrized by a suitable global complex variable $z$.
The moduli space of topological string theory is a special Kähler manifold. One of the consequences is that one can define a non-holomorphic object called the propagator $S(z,\bar{z})$ over the moduli space from the non-holomorphic genus one free energy by
$\partial_{z}F_{1}=\frac{1}{2}C(z)S+h(z).$ (3.1)
Here $C(z)$ is the Yukawa coupling, which is defined through the Griffiths transversality of the holomorphic (3,0) form $\Omega$ of the Calabi-Yau threefold, and which is holomorphic in $z$. $h(z)$ is also holomorphic in $z$, but it can be chosen by hand and regarded as a gauge choice of the theory.
The propagator $S$ has many nice properties. It satisfies a differential relation
$\partial_{z}S=C(z)\left(S^{2}+2\mathfrak{s}(z)S+\mathfrak{f}(z)\right),$ (3.2)
where both $\mathfrak{s}(z)$, $\mathfrak{f}(z)$ are holomorphic, therefore generating a differential ring called the BCOV ring Hosono:2008np . It is also related to the Levi-Civita connection $\Gamma^{z}_{zz}$ on the moduli space by
$\Gamma^{z}_{zz}=-C(z)(S+\mathfrak{s}(z)).$ (3.3)
Finally, the non-holomorphic free energies $F_{g}(z,\bar{z})$ of $g\geq 2$ can be written as polynomials in $S$ whose coefficients are rational functions in $z$, and they are subject to the BCOV holomorphic anomaly equations Bershadsky:1993cx ; Bershadsky:1993ta
$\partial_{S}F_{g}=\frac{1}{2}\left(\mathcal{D}^{2}F_{g-1}+\sum_{h=1}^{g-1}\mathcal{D}_{z}F_{h}\mathcal{D}_{z}F_{g-h}\right),\quad g\geq 2.$ (3.4)
Here $\mathcal{D}$ is the covariant derivative on the moduli space, acting on the free energies via
$\mathcal{D}_{z}F_{g}=\partial_{z}F_{g},\quad\mathcal{D}_{z}^{2}F_{g}=(\partial_{z}-\Gamma^{z}_{zz})\mathcal{D}_{z}F_{g}.$ (3.5)
Another important consequence of special Kähler manifold is that in each local open patch of the moduli space one can choose a symplectic basis of the period integrals $(T,T_{D})$ of the holomorphic (3,0) form $\Omega$, known as the choice of a frame $\Gamma$, so that $T$ is the flat coordinate on the local patch, and that the basis satisfies the special geometry relation
$T_{D}=c\frac{\partial\mathcal{F}_{0}(T)}{\partial T},$ (3.6)
where $\mathcal{F}_{0}$ is the prepotential or the genus zero free energy in the frame, and $c$ some normalization constant. Once a frame is chosen, one can reduce the propagator to its holomorphic limit
$S\rightarrow\mathcal{S}_{\Gamma}=-\frac{1}{C(z)}\frac{\partial_{z}^{2}T}{\partial_{z}T}-\mathfrak{s}(z),$ (3.7)
and likewise for the non-holomorphic free energies
$F_{g}(z,\bar{z})\rightarrow\mathcal{F}^{\Gamma}_{g}(z)=F_{g}(z,S=\mathcal{S}_{\Gamma}).$ (3.8)
We use Roman letters to denote non-holomorphic objects and caligraphic letters to denote the holomorphic limit.
In the case of chiral 2d Yang-Mills theory, by comparing (2.21) and (3.4), we can choose the global complex variable
z=t,z=-t, (3.9)
and $C(z)=1,\mathfrak{s}(z)=0$ so that $\Gamma^{z}_{zz}=-S$. The propagator $S$ given by (2.23) can fit into the definition (3.1) if we make the gauge choice $h(z)=-E_{2}(q)/24$. Finally, the propagator $S$ indeed enjoys the differential relation
$\partial_{z}S=-\partial_{t}S=S^{2},$ (3.10)
with simply $\mathfrak{s}(z)=\mathfrak{f}(z)=0$.
There are two important frames for the chiral 2d Yang-Mills. The first frame is the so-called “large radius” frame, where we choose $t$ as the flat coordinate, and the dual period integral is
$t_{D}=t^{2}/2,$ (3.11)
so that they satisfy the special geometry relation
$t_{D}=-\frac{\partial\mathcal{F}_{0}}{\partial t}.$ (3.12)
The period modulus $\tau$ defined by (2.9) can be written as
$\tau=\frac{{\mathsf{i}}}{2\pi}\frac{\partial t_{D}}{\partial t}=-\frac{{\mathsf{i}}}{2\pi}\frac{\partial^{2}\mathcal{F}_{0}}{\partial t^{2}}.$ (3.13)
Note that here the relationship between the flat coordinate $t$ and the global variable $z$ known as the mirror map is trivial, given by (3.9), unlike the usual case in topological string. The holomorphic free energies calculated in Section 2.2 are in fact the holomorphic limit of the non-holomorphic free energies in the large radius frame. Indeed, we find by (3.7) that in this limit
$S\rightarrow\mathcal{S}_{\text{LR}}=\frac{\partial_{t}^{2}t}{\partial_{t}t}=0,$ (3.14)
which can be equally obtained by sending $\bar{t}\rightarrow\infty$, and the non-holomorphic free energies $F_{g}(t,\bar{t})$ reduce to $\mathcal{F}_{g}(t)$.
The other important frame is the so-called “conifold” frame, where we exchange the roles of the two period integrals. We choose
$t_{c}=\frac{t^{2}}{2}$ (3.15)
as the flat coordinate and $t_{c,D}=t$ as the dual period integral. The conifold frame is related to the large radius frame by an $S$-transformation: this is seen clearly from $\tau$
$\tau_{{\rm c}}:=-2\pi{\mathsf{i}}\frac{\partial t_{cD}}{\partial t_{c}}=-\frac{1}{\tau}.$ (3.16)
The holomorphic limit of the propagator is
$S\rightarrow\mathcal{S}_{{\rm c}}=\frac{\partial_{t}^{2}t_{c}}{\partial_{t}t_{c}}=\frac{1}{t},$ (3.17)
which corresponds to sending $\bar{t}\rightarrow 0$. Equivalently, this limit can be obtained by performing the $S$-transformation (3.16) on the Eisenstein series $E_{2}(\tau),\widehat{E}_{2}(\tau)$ in (2.23), taking in account that
$E_{2}(\tau)=\tau^{-2}E_{2}(-1/\tau)+\frac{12}{t},\quad\widehat{E}_{2}(\tau)=\tau^{-2}\widehat{E}_{2}(-1/\tau),$ (3.18)
and then replacing $\widehat{E}_{2}$ back by $E_{2}$. Similarly, the holomorphic limit of the free energy in the conifold frame is obtained by performing the $S$-transform on all the Eisenstein series followed by the replacement $\widehat{E}_{2}\rightarrow E_{2}$ or by simply sending $\bar{t}\rightarrow 0$.
There are also two singular points in the moduli space. The first singular point is the “large radius” point located at
$t\rightarrow\infty,$ (3.19)
or $q={\rm e}^{-t}\rightarrow 0$, corresponding to the large $t$ regime of Okuyama:2019rqn . Near the large radius point, the holomorphic limit of the free energy $\mathcal{F}_{g}$ in the large radius frame has the asymptotic behavior
$\mathcal{F}^{\text{LR}}_{g}=\mathcal{F}_{g}\sim\frac{q^{2}}{(g-1)(2g-3)!}+\mathcal{O}(q^{3}),\quad g\geq 2.$ (3.20)
The second singular point is the “conifold” point located at
$t=0,$ (3.21)
and it corresponds to the small $t$ regime of Okuyama:2019rqn . Near the conifold point, the holomorphic limit of the free energy $\mathcal{F}_{g}^{{\rm c}}$ in the conifold frame has the asymptotic behavior [^7] If we choose the naive non-holomorphic uplift of $\mathcal{F}_{g}$, simply replacing all $E_{2}$ by $\widehat{E}_{2}$, the conifold frame free energy $\mathcal{F}^{{\rm c}}_{g}$ would then not display gapped asymptotic behavior near $t=0$. This is another piece of evidence that the naive non-holomorphic uplift of free energy does not work. Okuyama:2019rqn
$\mathcal{F}^{{\rm c}}_{g}\sim\frac{(-1)^{g}16\sqrt{2}\pi^{2g}(4g-5)!B_{2g}(1/2)}{2^{2g}(2g)!(2g-3)!t_{c}^{2g-3/2}}+\mathcal{O}({\rm e}^{-4\pi^{2}/t}),\quad g\geq 2,$ (3.22)
This is similar to the gap condition of the conifold frame free energy of topological string theory near a conifold singular point Ghoshal:1995wm , although the leading exponent of the flat coordinate and the leading coefficient are different.
4 Instanton amplitudes and resurgent structure
4.1 General strategy
We study in the following non-perturbative corrections to the free energy of the topological string theory dual to chiral 2d Yang-Mills. We first describe our general strategy.
There could be various instanton sectors with different instanton actions $\mathcal{A}_{\alpha}$ whose contributions to the non-perturbative corrections, also called the instanton amplitudes, are of the order ${\rm e}^{-\mathcal{A}_{\alpha}/g_{s}}$. Here we demand that the different instanton sectors are such that their instanton actions as complex numbers are not colinear, i.e.
$\arg\mathcal{A}_{\alpha}-\arg\mathcal{A}_{\beta}\not\in 2\pi\mathbb{Z},\quad\forall\alpha\neq\beta.$ (4.1)
In addition, in an instanton sector with action $\mathcal{A}$, there could be multi-instanton amplitudes $F^{(n)}$ of the order ${\rm e}^{-n\mathcal{A}/g_{s}}$ with $n\geq 1$ so that the complete non-perturbative correction from this instanton sector takes the form of a trans-series,
$F^{\text{np}}=F^{(0)}+\sum_{n=1}^{\infty}F^{(n)}.$ (4.2)
Here $F^{(0)}$ is the power series of the perturbative free energy
$F^{(0)}=\sum_{g=0}^{\infty}g_{s}^{2g-2}F_{g},$ (4.3)
and the instanton amplitude $F^{(n)}$ at order $n$ takes the form
$F^{(n)}={\rm e}^{-n\mathcal{A}/g_{s}}g_{s}^{\nu_{n}}\sum_{k=0}^{\infty}g_{s}^{k}F_{k}^{(n)}.$ (4.4)
Our general strategy to find instanton amplitudes is based on two crucial assumptions following Couso-Santamaria2016 ; Couso-Santamaria2015 ; Gu:2022sqc ; Gu:2023mgf . The first assumption is that the trans-series (4.2) including all instanton contributions from a particular instanton sector also satisfies the BCOV holomorphic anomaly equations. We can then expand the holomorphic anomaly equations in terms of two small parameters $g_{s}$ and ${\rm e}^{-\mathcal{A}/g_{s}}$ leading to recursion relations for $F_{k}^{(n)}$, which can be solved by direct integration Couso-Santamaria2016 ; Couso-Santamaria2015 . In this process, just like solving the perturbative free energy $F_{g}$, there will also be holomorphic ambiguities in the instanton amplitudes. To fix them, we make the second assumption that the instanton amplitudes are closely related to the perturbative free energy in accord with the expectation of the resurgence theory Ecalle . What this second assumption means is the following.
The trans-series (4.2) is a formal object as both the perturbative power series $F^{(0)}$ and the power series in each instanton amplitude $F^{(n)}$ are divergent power series of the 1-Gevrey type, whose coefficients grow factorially fast,
$F_{g}\sim(2g-2)!,\quad F_{k}^{(n)}\sim k!.$ (4.5)
In general given a divergent series of 1-Gevrey type
$\varphi(z)=\sum_{n=0}^{\infty}a_{n}z^{n},\quad a_{n}\sim n!,$ (4.6)
we can resum it and convert it to a finite number by the means of Borel resummation. The Borel resummation is the Laplace transform of the Borel transform
$\mathscr{S}\varphi(z)=\frac{1}{z}\int_{0}^{{\rm e}^{{\mathsf{i}}\arg z}\infty}\mathcal{B}[\varphi](\zeta){\rm e}^{-\zeta/z}{\rm d}\zeta,$ (4.7)
which in turn is a convergent power series defined by
$\mathcal{B}[\varphi](\zeta)=\sum_{n=0}^{\infty}\frac{a_{n}}{n!}\zeta^{n}.$ (4.8)
The Borel resummation is well-defined if the Borel transform as the sum of a convergent series can be analytically continued along the direction of $\arg z$ to infinity. This is possible if along this direction there are no singular points of the Borel transform, also known as Borel singularities. For each Borel singularity $\mathcal{A}_{\alpha}$ with $\vartheta_{\alpha}=\arg\mathcal{A}_{\alpha}$, define a Stokes ray $\rho_{\alpha}={\rm e}^{{\mathsf{i}}\vartheta_{\alpha}}\mathbb{R}_{+}$ in the complex $z$ plane. Then the Borel resummation defines a continuous function in disjoint cones in the complex $z$ plane, which are separated by Stokes rays. The discontinuity of the function across a Stokes ray $\rho_{\alpha}$ is given by the discrepancy of the pair of lateral Borel resummations, whose integration contours sandwich the Stokes ray, i.e.
$\text{disc}_{\vartheta_{\alpha}}\varphi(z)=\mathscr{S}^{(+)}\varphi(z)-\mathscr{S}^{(-)}\varphi(z),$ (4.9)
with
$\mathscr{S}^{(\pm)}\varphi(z)=\frac{1}{z}\int_{0}^{{\rm e}^{{\mathsf{i}}\vartheta_{\alpha}\pm{\mathsf{i}}0^{\pm}}\infty}\mathcal{B}[\varphi](\zeta){\rm e}^{-\zeta/z}{\rm d}\zeta.$ (4.10)
This is known as the Stokes discontinuity, and it is of the order of magnitude ${\rm e}^{-\mathcal{A}_{\alpha}/z}$, indicative of possible non-perturbative corrections. In fact, by the resurgence theory, the Borel singularity $\mathcal{A}_{\alpha}$ that gives rise to the Stokes ray $\rho_{\alpha}$ is the action of an instanton sector. In the simple case where there are no multi-instantons (only one Borel singularity on the Stokes ray), the Stokes discontinuity is simply given by the Borel resummation of the associated instanton amplitude
$\text{disc}_{\vartheta_{\alpha}}\varphi(z)=\mathcal{S}_{\alpha}\mathscr{S}^{(-)}\varphi^{(\alpha)}(z),$ (4.11)
with
$\varphi^{(\alpha)}(z)={\rm e}^{-\mathcal{A}_{\alpha}/z}z^{-\nu_{\alpha}}\sum_{k=0}^{\infty}a_{k}^{(\alpha)}z^{k},$ (4.12)
up to a proportionality constant $\mathcal{S}_{\alpha}$ known as the Stokes constant. The formula of Stokes discontinuity (4.11) also implies the famous resurgence relation: the coefficients of the instanton amplitudes control the asymptotic behavior of the perturbative coefficients
$a_{n}\sim\frac{\mathcal{S}_{\alpha}}{2\pi{\mathsf{i}}}\frac{\Gamma(n+\nu_{\alpha})}{\mathcal{A}^{n+\nu_{\alpha}}}\sum_{k=0}^{\infty}\frac{a_{k}^{(\alpha)}\mathcal{A}_{\alpha}^{k}}{\prod_{j=1}^{k}(n+\nu_{\alpha}-j)}.$ (4.13)
The formula of Stokes discontinuity (4.11) and resurgence relation (4.13) allow us to extract information of the instanton amplitudes from the perturbative power series alone.
If there are different instanton sectors, their contributions add up in the resurgence relation. Obviously, the instanton sector whose instanton action has the smallest absolute value plays the dominant role. On the other hand, in the case of resonant resurgence where instanton sectors arise in pairs with opposite instanton actions, their contributions to odd power perturbative coefficients cancel, and one gets
$\varphi(z)=\sum_{n=0}^{\infty}a^{\prime}_{n}z^{2n},\quad a^{\prime}_{n}\sim(2n)!,$ (4.14)
whose remaining coefficients satisfy the resurgence relation
$a^{\prime}_{n}\sim\frac{\mathcal{S}_{\alpha}}{\pi{\mathsf{i}}}\frac{\Gamma(2n+\nu_{\alpha})}{\mathcal{A}_{\alpha}^{2n+\nu_{\alpha}}}\sum_{k=0}^{\infty}\frac{a_{k}^{(\alpha)}\mathcal{A}_{\alpha}^{k}}{\prod_{j=1}^{k}(2n+\nu_{\alpha}-j)}.$ (4.15)
This is what we will encounter in topological string.
For generic cases where there are multi-instanton contributions with actions $\ell\mathcal{A}$ $(\ell=2,3,\ldots)$, the impact of the higher order instanton amplitudes is more complicated, although it can still be described in a precise way as discussed in more detail in Section 4.4. In the special case where the multi-instanton amplitudes truncate to finite powers, their contributions to the formula of Stokes discontinuity also simply add up, just like contributions from different instanton sectors.
The instanton sectors which contribute to the Stokes discontinuity of the perturbative power series are said to form the resurgent structure from the perturbative series Gu:2021ize . They may constitute only a subset of all possible non-perturbative contributions, but they are the most important ones. Here, we will only be interested in this type of non-perturbative contributions to the topological string theory dual to chiral 2d Yang-Mills.
These non-perturbative contributions to topological string theory have been studied in other examples, for instance when the background geometry is a local Calabi-Yau threefold Gu:2022sqc or a simple compact Calabi-Yau threefold Gu:2023mgf (see Marino:2023nem ; Alexandrov:2023wdj ; Gu:2024wag also for generalization to real topological string and refined topological string as well as to other observables in topological string.). Several observations were made which were then conjectured to be true for generic Calabi-Yau background. We will test these conjectures for the background geometry $X_{E}$, the double line bundle over an elliptic curve (2.5).
It was demonstrated in Gu:2022sqc ; Gu:2023mgf that the resurgent assumption puts very strong constraints on possible instanton sectors. The instanton actions are in fact all integral periods of the Calabi-Yau background Drukker:2011zy ; Couso-Santamaria2016 ; Couso-Santamaria2015 , each of which can be used to define the flat coordinate of a symplectic frame, known as the $A$-frame of the instanton sector. In addition, given an instanton sector with action $\mathcal{A}$, the holomorphic limit of the instanton amplitudes in the $A$-frame simplify greatly, reducing to truncated power series of finite degrees, and they can be entirely determined by the resurgent properties of the perturbative free energy. In turn, these $A$-frame instanton amplitudes can serve as boundary conditions to fix completely the holomorphic ambiguity of the instanton amplitudes as solutions to the holomorphic anomaly equations.
(a) $t=4\pi$
(b) $t=\pi$
Figure 4.1: Borel singularities of perturbative free energy in the large radius frame. We use perturbative series truncated to 200 terms, and use Padé approximant to mimic the analytic continuation of the Borel transform. The singular points (red) of the approximation would condense to branch cuts if the truncation is pushed to infinity. At $t=4\pi$ (a), the branch points (black) have charges $\pm(1,0,0),\pm(1,2,2),\pm(1,-2,2)$. At $t=\pi$ (b), the branch points (black) have charges $\pm(1,0,0)$.
We show in this section that all these observations still hold true for the topological string theory dual to the chiral 2d Yang-Mills, although there are subtle differences. First of all, simple plots of Borel singularities of perturbative free energy as in Figs. 4.1 show that the instanton actions are periods of the form
$\mathcal{A}_{\gamma}=\alpha t^{2}/2+\beta 2\pi{\mathsf{i}}\,t+\delta\pi^{2},$ (4.16)
with the charge vector $\gamma=(\alpha,\beta,\delta)\in\mathbb{Z}^{3}$. The dominant instanton sectors have actions
$\pm\mathcal{A}_{c}=\pm t^{2}/2,$ (4.17)
and we call them real instantons as the actions are real for $t>0$. In Section 4.2 we solve completely both the 1-instanton and multi-instanton amplitudes in the real instanton sectors in the $A$-frame by exploiting the resurgence relations and the formula of Stokes discontinuity. These formulas are valid even when $\theta$ is turned on so that $t$ is complex. Using these as boundary conditions, we fix the holomorphic ambiguities of the trans-series solutions to the holomorphic anomaly equations, first for 1-instanton amplitude in Section 4.3, and then for multi-instanton amplitudes in Section 4.4. We check these solutions by taking them to non-$A$-frames and comparing with numerical calculation of the Stokes discontinuity of the perturbative free energy.
Based on these results, we write down in Section 4.5 the non-perturbative topological string partition function which takes into account non-perturbative corrections from all the real instantons and which is real for positive $t$ and $g_{s}$. There is in fact an infinite family of them, and the simplest member of it is the medium Borel resummation of the perturbative partition function. We also compare with the proposal of Okuyama and Sakai Okuyama:2018clk summarized in Section 2.3.
Finally, in Section 4.6 we study the spectrum of complex instanton sectors with actions different from (4.17). We find two infinite towers of complex instantons as well as a pair of additional complex instantons for which the 1-instanton amplitude solved in Section 4.3 works universally. We expect they correspond to BPS bound states of D-branes in type II string. Their wall-crossing behavior is also discussed.
4.2 Boundary conditions
We first explore instanton amplitudes for the real instantons in the $A$-frame, which is nothing else but the conifold frame. Recall that the the free energy $\mathcal{F}_{g}^{{\rm c}}$ in the conifold frame displays the gap-like asymptotic behavior (3.22). In the leading order, it reads
$\mathcal{F}_{g}^{{\rm c}}\sim\frac{\Gamma(2g-\frac{3}{2})}{\mathcal{A}_{c}^{2g-\frac{3}{2}}}\sqrt{\frac{2}{\pi}}\left(1+\mathcal{O}(1/2)^{2g}\right),\quad g\geq 2,$ (4.18)
with
$\mathcal{A}_{c}=t_{c}=t^{2}/2.$ (4.19)
By comparing wth the resurgence relation (4.15), one concludes that the dominant instanton sectors are indeed the pair with instanton actions $\pm\mathcal{A}_{c}$ Okuyama:2018clk . Furthermore, in the $A$-frame, denoted by the subscript $\mathcal{A}$, the 1-instanton amplitude is simply [^8] The prefactor $g_{s}^{-2}$ is to match the leading exponent of the perturbative free energy.,
$\mathcal{F}^{(1),{\rm c}}_{\mathcal{A}}=g_{s}^{-2}\pi{\mathsf{i}}\sqrt{\frac{2}{\pi}}g_{s}^{3/2}{\rm e}^{-\mathcal{A}_{c}/g_{s}},$ (4.20)
which has only one term in the power series. Note that it is different from the usual case in topological string where 1-instanton amplitudes truncate to two terms in the $A$-frame Pasquetti2010 .
On the other hand, the gap-like asymptotic behavior (3.22) has subleading contributions other than (4.18), which implies possible multi-instanton amplitudes. To uncover these higher order instanton amlitudes, we calculate the Borel resummation of the following divergent power series
$\mathcal{F}^{\text{gap}}(g_{s})=\sum_{g=2}^{\infty}a_{g}g_{s}^{2g-2},$ (4.21)
that describes the gap condition, where $a_{g}$ are the coefficients of the gap condition (3.22), given by
$a_{g}=\frac{(-1)^{g}16\sqrt{2}\pi^{2g}(4g-5)!B_{2g}(1/2)}{2^{2g}(2g)!(2g-3)!}.$ (4.22)
Let $z=g_{s}/t_{c}=g_{s}/\mathcal{A}_{c}$, the divergent power series reads
$\mathcal{F}^{\text{gap}}(g_{s})=$ $\frac{g_{s}^{1/2}}{t_{c}}\sum_{g=2}^{\infty}a_{g}z^{2g-\frac{5}{2}}$
$=$ $\frac{g_{s}^{1/2}}{t_{c}}\sum_{g=2}^{\infty}\frac{a_{g}}{(2g-\frac{5}{2})!}z^{2g-\frac{5}{2}}\int_{0}^{\infty}{\rm e}^{-\zeta}\zeta^{2g-\frac{5}{2}}{\rm d}\zeta$
$\sim$ $\frac{g_{s}^{1/2}}{t_{c}}\frac{1}{z}\int_{0}^{{\rm e}^{{\mathsf{i}}\vartheta}\infty}{\rm e}^{-\zeta/z}\sum_{g=2}^{\infty}\frac{a_{g}}{(2g-\frac{5}{2})!}\zeta^{2g-\frac{5}{2}}{\rm d}\zeta,\quad\vartheta=\arg z.$ (4.23)
Here in the third step we illegally exchange the orders of integration and summation, which is the reason that we obtain in the end a convergent integral, which is nothing else but the Borel resummation of the original divergent power series. In other words, we find
$\mathscr{S}\mathcal{F}^{\text{gap}}(g_{s})=g_{s}^{-1/2}\int_{0}^{{\rm e}^{{\mathsf{i}}\vartheta}\infty}{\rm e}^{-\zeta\mathcal{A}_{c}/g_{s}}\sum_{g=2}^{\infty}\frac{a_{g}}{(2g-\frac{5}{2})!}\zeta^{2g-\frac{5}{2}}{\rm d}\zeta.$ (4.24)
To evaluate the Borel resummation, we use the fact that
$B_{2g}(1/2)=(-1)^{g}\frac{2(2g)!}{(4\pi)^{2g}}\sum_{\ell=0}^{\infty}\left(\frac{1}{(\ell+1/2)^{2g}}-\frac{1}{(\ell+1)^{2g}}\right).$ (4.25)
This allows us to derive the following key identity
$\sum_{g=2}^{\infty}\frac{a_{g}}{(2g-\frac{5}{2})!}\zeta^{2g-\frac{5}{2}}=-\sqrt{\frac{2}{\pi}}\left(\frac{\pi^{2}}{12\zeta^{1/2}}+\frac{1}{2\zeta^{5/2}}-\frac{\pi}{2\zeta^{3/2}\sin(\pi\zeta)}\right),$ (4.26)
which gives the analytic continuation of the Borel transform inside the formula of Borel resummation (4.24). The latter can then be written as
$\mathscr{S}\mathcal{F}^{\text{gap}}(g_{s})=-\sqrt{\frac{2}{\pi g_{s}}}\int_{0}^{{\rm e}^{{\mathsf{i}}\vartheta}\infty}{\rm e}^{-\zeta\mathcal{A}_{c}/g_{s}}\left(\frac{\pi^{2}}{12\zeta^{1/2}}+\frac{1}{2\zeta^{5/2}}-\frac{\pi}{2\zeta^{3/2}\sin(\pi\zeta)}\right){\rm d}\zeta.$ (4.27)
As usually happens in Borel resummation, the integral above is obstructed when the integrand has singular points along the integration contour. This happens when
$\vartheta=0\;\Leftrightarrow\;\arg g_{s}=\arg\mathcal{A}_{c},$ (4.28)
and the integration is along the positive real axis. The Stokes discontinuity is defined by
$\text{disc}_{0}\mathcal{F}^{\text{gap}}(g_{s})=-\sqrt{\frac{2}{\pi g_{s}}}\int_{\mathcal{H}}{\rm e}^{-\zeta\mathcal{A}_{c}/g_{s}}\left(\frac{\pi^{2}}{12\zeta^{1/2}}+\frac{1}{2\zeta^{5/2}}-\frac{\pi}{2\zeta^{3/2}\sin(\pi\zeta)}\right){\rm d}\zeta,$ (4.29)
where $\mathcal{H}$ is the Hankel type integration contour that comes from $\infty$ below the positive real axis, half-circles the origin, and goes away to $\infty$ above the positive real axis. The integrand has a string of poles at [^9] Note that the integrand is not singular at the origin, as it is the analytic continuation of the Borel transform of $\mathcal{F}^{\text{gap}}(g_{s})$, which is a convergent series with a positive radius of convergence.
$\zeta=n,\quad n=1,2,3,\ldots$ (4.30)
on the positive real axis. The Stokes discontinuity is then evaluated by summing up residues of all these poles, and we arrive at
$\text{disc}_{0}\mathcal{F}^{\text{gap}}(g_{s})={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}\sum_{n=1}^{\infty}\frac{(-1)^{n-1}}{n^{3/2}}{\rm e}^{-n\mathcal{A}_{c}/g_{s}}.$ (4.31)
This indicates that indeed there are multi-instanton contributions whose actions are
$n\mathcal{A}_{c}=nt_{c}=\frac{nt^{2}}{2},$ (4.32)
and whose instanton amplitudes in the $A$-frame are
$\mathcal{F}^{(n)}_{\mathcal{A}_{c}}=\mu_{n}{\rm e}^{-n\mathcal{A}_{c}/g_{s}},\quad\mu_{n}={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{n-1}}{n^{3/2}},$ (4.33)
all of which truncate to a single term, in accord with our expectation. Note again that they are slightly different from the usual case in topological string, where in the $A$-frame all multi-instanton amplitudes truncate to two terms Gu:2022sqc .
4.3 1-instanton amplitudes
To find the instanton amplitudes for the $\pm\mathcal{A}_{c}$ sectors in a non-$A$-frame, we solve the trans-series solutions to the holomorphic anomaly equations or HAEs (2.21). We solve 1-instanton amplitude in this subsection and multi-instanton amplitudes in the next. In this and the next subsections, we drop the subscript $c$ in the action $\mathcal{A}$, since the solutions we find are universal and apply for other instanton sectors as well, as we will see in Section 4.6.
Let us define
$\widetilde{F}=\sum_{g\geq 1}g_{s}^{2g}F_{g},\quad\widehat{F}=\sum_{g\geq 2}g_{s}^{2g}F_{g}.$ (4.34)
The HAEs (2.21) can be written in the master form
$\partial_{S}\widehat{F}=\frac{1}{2}g_{s}^{2}\mathcal{D}^{2}\widetilde{F}+\frac{1}{2}(\mathcal{D}\widetilde{F})^{2}.$ (4.35)
By our first assumption, the trans-series (4.2) should also satisfy the HAE (4.35) with
$\widehat{F}^{(\ell)}=\widetilde{F}^{(\ell)}=g_{s}^{2}F^{(\ell)},\quad\ell\geq 1.$ (4.36)
Expanding up to the 1-instanton order, we find the following linear equation for $F^{(1)}$,
$\partial_{S}F^{(1)}=\frac{g_{s}^{2}}{2}\mathcal{D}^{2}F^{(1)}+\mathcal{D}\widetilde{F}^{(0)}\mathcal{D}F^{(1)}.$ (4.37)
It was found in Gu:2022sqc that
$\text{const.}\exp\left(-\Phi^{(n)}/g_{s}\right)$ (4.38)
satisfies the linear equation (4.37). Here the function $\Phi^{(n)}$ in the exponent is,
$\Phi^{(n)}=\frac{1}{g_{s}\mathbb{D}}\left(1-{\rm e}^{-ng_{s}\mathbb{D}}\right)G=\sum_{k=1}^{\infty}\frac{(-g_{s})^{k-1}n^{k}}{k!}\mathbb{D}^{k-1}G$ (4.39)
with the derivative $\mathbb{D}$,
$\mathbb{D}=\mathcal{D}\mathcal{A}(S-\mathcal{S}_{\mathcal{A}})\mathcal{D},$ (4.40)
and $\mathcal{S}_{\mathcal{A}}$ the holomorphic propagotor in the $A$-frame. The function $G$ is
$G=\mathcal{A}+\mathbb{D}\widetilde{F}^{(0)}.$ (4.41)
To fix the normalization constant and the exponent $n$, we evaluate the instanton amplitude in the holomorphic limit of the $A$-frame. This is realized by
$S\rightarrow\mathcal{S}_{\mathcal{A}},$ (4.42)
where the derivative $\mathbb{D}$ simply vanishes, and the expoential in (4.38) reduces to ${\rm e}^{-n\mathcal{A}/g_{s}}$. By comparing with (4.20), we can fix the normalization constant and the exponent, and the 1-instanton amplitude is
$F^{(1)}={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}\exp\left(-\Phi^{(1)}/g_{s}\right).$ (4.43)
We can also work out the holomorphic limit of the 1-instanton amplitude in a non-$A$-frame. As argued in Gu:2022sqc and demonstrated below, the propagator $S$ has the property that when replaced by its holomorphic limit $\mathcal{S}$ in a non-$A$-frame with flat coordinate $T$, its difference from $\mathcal{S}_{\mathcal{A}}$ can always be written as
$\mathcal{S}-\mathcal{S}_{\mathcal{A}}=\frac{\alpha}{\mathcal{D}T\mathcal{D}\mathcal{A}},$ (4.44)
so that
$\mathbb{D}\rightarrow\alpha\partial_{T}.$ (4.45)
Here $\alpha$ is a frame dependent constant and in fact is the coefficient of the dual period in the decomposition of $\mathcal{A}$ in terms of the symplectic basis
$\mathcal{A}=\alpha T_{D}+\ldots$ (4.46)
When $\alpha\neq 0$, we can then always redefine the prepotential $\mathcal{F}_{0}$ so that
$\mathcal{A}=\alpha\partial_{T}\mathcal{F}_{0}.$ (4.47)
We find then that in this limit,
$G\rightarrow g_{s}^{2}\alpha\partial_{T}\mathcal{F}^{(0)},$ (4.48)
and
$\Phi^{(1)}\rightarrow g_{s}\left(\mathcal{F}^{(0)}(T)-\mathcal{F}^{(0)}(T-\alpha g_{s})\right).$ (4.49)
The holomorphic limit of the 1-instanton amplitude $F^{(1)}$ is then  [^10] This solution can also be obtained directly from (4.20) via a frame transformation Marino:2024yme ; Marino:2024tbx .
$F^{(1)}\rightarrow\mathcal{F}^{(1)}={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}\exp\left(\mathcal{F}^{(0)}(T-\alpha g_{s})-\mathcal{F}^{(0)}(T)\right).$ (4.50)
More explicitly, when expanded in small $g_{s}$, we have the trans-series form
$\mathcal{F}^{(1)}={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}{\rm e}^{-\mathcal{A}/g_{s}}{\rm e}^{\frac{1}{2}\alpha^{2}t+\pi{\mathsf{i}}\alpha\beta}\exp\left(-\frac{g_{s}\alpha^{3}}{6}+\sum_{n=1}g_{s}^{n}\sum_{g=1}^{\lfloor\frac{n+1}{2}\rfloor}\frac{(-\alpha)^{n+2-2g}}{(n+2-2g)!}\partial_{T}^{n+2-2g}\mathcal{F}_{g}(T)\right).$ (4.51)
If we evaluate the 1-instanton amplitude for the dominant instanton sector with action $\mathcal{A}_{c}$ in the large radius frame, we find from (3.14),(3.17) that (4.44) is indeed true with $\alpha=-1$, and the 1-instanton amplitude (4.51) reads
$\mathcal{F}^{(1)}={\mathsf{i}}\sqrt{\frac{2\pi}{g_{s}}}\exp\left(\mathcal{F}^{(0)}(t+g_{s})-\mathcal{F}^{(0)}(t)\right).$ (4.52)
This agrees with the 1-instanton amplitude from Okuyama:2018clk summarized in (2.41).
4.4 Multi-instanton amplitudes
In this section, we solve multi-instanton contributions to the non-perturbative free energy. For reasons that will become clear, it is more convenient to consider instead the non-perturbative partition function in an instanton sector,
$Z=Z^{(0)}+Z^{(1)}+\ldots,$ (4.53)
or rather the reduced partition function,
$Z_{r}=\frac{Z}{Z^{(0)}}=1+\sum_{n\geq 1}Z_{r}^{(n)},$ (4.54)
which is related to the non-perturbative free energy in the same instanton sector by
$Z_{r}=\exp\Bigg(\sum_{n\geq 1}F^{(n)}\Bigg).$ (4.55)
In the $A$-frame, the non-perturbative free energy reduces to a linear combination of ${\rm e}^{-n\mathcal{A}/g_{s}}$ with constant coefficients, and so should the reduced partition function,
$\mathcal{Z}_{r,\mathcal{A}}=\sum_{n=1}^{\infty}c_{n}{\rm e}^{-n\mathcal{A}/g_{s}}.$ (4.56)
We will use these as boundary conditions to fix the reduced partition function as solutions to the holomorphic anomaly equations.
Assuming that both $\log Z^{(0)}$ and $\log Z$ satisfy the holomorphic anomaly equation (4.35), one can derive that the reduced partition function satisfies the following version of holomorphic anomaly equation
$\partial_{S}Z_{r}=\frac{g_{s}^{2}}{2}\mathcal{D}^{2}Z_{r}+\mathcal{D}\widetilde{F}^{(0)}\mathcal{D}Z_{r}.$ (4.57)
It is similar to (4.37) and is also linear and homogeneous. This means that we can focus on the simplest boundary condition
$\mathcal{Z}_{r,\mathcal{A}}^{(n)}=c_{n}{\rm e}^{-n\mathcal{A}/g_{s}},$ (4.58)
with constant $c_{n}$ and look for the corresponding solution. In addition, given the similarity with (4.37), we find the same solution as (4.38),
$Z_{r}^{(n)}=c_{n}\exp(-\Phi^{(n)}/g_{s}),$ (4.59)
with now the exponent $n$ matching the instanton order and the specific normalization constant $n$. In the holomorphic limit of the $A$-frame, it is easy to see that (4.59) indeed reduces to (4.58). On the other hand, in the holomorphic limit of a non-$A$-frame, the derivative $\mathbb{D}$ becomes
$\mathbb{D}=\alpha\partial_{T}$ (4.60)
and we find that
$\mathcal{Z}_{r}^{(n)}=c_{n}{\rm e}^{\mathcal{F}(T-n\alpha g_{s})-\mathcal{F}(T)}=\frac{1}{Z^{(0)}}c_{n}{\rm e}^{-n\alpha g_{s}\partial_{T}}Z^{(0)}.$ (4.61)
This is the proposition that we will use repeatedly in the following: once its boundary condition in the $A$-frame is known, the reduced partition function in a non-$A$-frame can be obtained by the rule of replacelement
${\rm e}^{-\mathcal{A}/g_{s}}\rightarrow{\rm e}^{-\alpha g_{s}\partial_{T}}.$ (4.62)
Let us now clarify what kind of reduced partition function we wish to calculate, in other words, what exactly are the appropriate boundary conditions, and what are the constants $c_{n}$ in (4.56). Recall that we wish to focus on the non-perturbative corrections in the resurgent structure of the perturbative series. In particular, they should appear on the right hand side of the formula of Stokes discontinuity (4.11) across certain Stokes ray. It turns out that in the case that there are multi-instanton amplitudes, instead of Stokes discontinuity, it is more useful to consider the Stokes transformation $\mathfrak{S}_{\vartheta_{\alpha}}$ associated to the Stokes ray $\rho_{\alpha}$ Ecalle . It is an automorphism of trans-series,
$\mathfrak{S}_{\vartheta_{\alpha}}:\;\varphi(z)\rightarrow\varphi(z)+\mathcal{S}_{\alpha}\varphi^{(\alpha)}(z)+\ldots$ (4.63)
so that,
$\mathscr{S}^{(+)}\varphi(z)=\mathscr{S}^{(-)}\mathfrak{S}_{\vartheta_{\alpha}}\varphi(z).$ (4.64)
As an automorphism, the Stokes transformation can be written as exponential of differential operators known as pointed alien derivatives Ecalle
$\mathfrak{S}=\exp\left(\sum_{n=1}^{\infty}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\right).$ (4.65)
The alien derivative $\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}$ serves to send a trans-series of the order ${\rm e}^{-\ell\mathcal{A}/g_{s}}$ to another trans-series of the order ${\rm e}^{-(\ell+n)\mathcal{A}/g_{s}}$, and it behaves truely like a derivative, satisfying the Leibniz rule and the chain rule. In terms of alien derivatives, the action of the Stokes transformation reads
$\mathfrak{S}\varphi(z)=\varphi(z)+\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}\varphi(z)+\left(\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}+\frac{1}{2}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}\right)\varphi(z)+\ldots$ (4.66)
This formula tells us that when there are multi-instantons, their amplitudes are in general not unique [^11] This has also been emphasized in the trans-series formalism of eigen-energies in quantum mechanics, where there are $n$ instanton amplitudes at the $n$-instanton order. See for instance vanSpaendonck:2023znn .. At the order of ${\rm e}^{-2\mathcal{A}/g_{s}}$, there are two instanton amplitudes
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}\varphi(z),\;\;\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}\varphi(z),$ (4.67)
and at the order of ${\rm e}^{-3\mathcal{A}/g_{s}}$, there are three instanton amplitudes
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{3\mathcal{A}}\varphi(z),\;\;\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}\varphi(z),\;\;\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 3}}\varphi(z),$ (4.68)
and etc. Only in the special case where $\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\varphi(z)$ truncate to polynomials of finite degrees so that further application of alien derivatives vanishes do we have a single instanton amplitude at each order.
In the case of topological string free energy, we will first work out the basic $n$-instanton amplitude $F^{(n)}$ defined by,
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}F^{(0)}=F^{(n)}.$ (4.69)
Here we have absorbed the Stokes constant inside $F^{(n)}$ or simply set it to one. All the other $n$-instanton amplitudes can be derived from the basic instanton amplitudes as we will see later.
We first work out the boundary conditions, i.e. the instanton amplitudes in the $A$-frame. On the one hand, we know from (4.33) that
$\mathfrak{S}\mathcal{F}^{(0)}_{\mathcal{A}}=\mathcal{F}^{(0)}+\sum_{n=1}^{\infty}\mu_{n}{\rm e}^{-n\mathcal{A}/g_{s}}.$ (4.70)
On the other hand, (4.65) implies
$\mathfrak{S}\mathcal{F}^{(0)}_{\mathcal{A}}=\mathcal{F}^{(0)}+\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}\mathcal{F}^{(0)}_{\mathcal{A}}+\left(\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}+\frac{1}{2}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}\right)\mathcal{F}^{(0)}_{\mathcal{A}}$$\ldots$ (4.71)
Given that the alien derivatives annihilate truncated power series, we find recursively that
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\mathcal{F}^{(0)}_{\mathcal{A}}=\mathcal{F}^{(n)}_{\mathcal{A}}=\mu_{n}{\rm e}^{-n\mathcal{A}/g_{s}}.$ (4.72)
To find the basic instanton amplitude $F^{(n)}$ at order $n$, we single out the contribution of $\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}$ and ignore for the moment all the other alien derivatives. We choose the special boundary condition that
$\Gamma_{n}:\quad\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\ell\mathcal{A}}\mathcal{F}^{(0)}_{\mathcal{A}}=\delta_{n,\ell}\mathcal{F}^{(n)}_{\mathcal{A}}$, (4.73)
so that
$\Gamma_{n}:\quad\mathfrak{S}\mathcal{F}^{(0)}_{\mathcal{A}}=\mathcal{F}^{(0)}_{\mathcal{A}}+\mathcal{F}^{(n)}_{\mathcal{A}}$. (4.74)
We use the label $\Gamma_{n}$ to emphasize the special boundary condition that we have chosen. The partition function then reads
$\Gamma_{n}:\quad\mathfrak{S}\mathcal{Z}^{(0)}_{\mathcal{A}}=\mathcal{Z}_{\mathcal{A}}^{(0)}\exp\mathcal{F}_{\mathcal{A}}^{(n)},$ (4.75)
and the reduced partition function after crossing the Stokes ray is
$\Gamma_{n}:\quad\mathcal{Z}_{r,\mathcal{A}}=\exp\mathcal{F}^{(n)}_{\mathcal{A}}=\sum_{\ell=0}^{\infty}\frac{1}{\ell!}\mu_{n}^{\ell}{\rm e}^{-n\ell\mathcal{A}/g_{s}}.$ (4.76)
By our replacement rule (4.62), the reduced partition function in a non-$A$-frame is
$\Gamma_{n}:\quad\mathcal{Z}_{r}=\sum_{\ell=0}^{\infty}\frac{1}{\ell!}\mu_{n}^{\ell}{\rm e}^{\mathcal{F}^{(0)}(T-n\ell\alpha g_{s})-\mathcal{F}^{(0)}(T)},$ (4.77)
and the non-perturbative free energy is
$\Gamma_{n}:\quad\mathcal{F}=\log\mathcal{Z}=\mathcal{F}^{(0)}+\mu_{n}{\rm e}^{\mathcal{F}^{(0)}(T-n\alpha g_{s})-\mathcal{F}^{(0)}(T)}+\mathcal{O}({\rm e}^{-2n\mathcal{A}/g_{s}}).$ (4.78)
By comparing with
$\Gamma_{n}:\quad\mathcal{F}=\mathfrak{S}\mathcal{F}^{(0)}=\left(\exp\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\right)\mathcal{F}^{(0)}=\mathcal{F}^{(0)}+\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\mathcal{F}^{(0)}+\mathcal{O}({\rm e}^{-2n\mathcal{A}/g_{s}}),$ (4.79)
we conclude that the basic instanton amplitude at order $n$ is
$\mathcal{F}^{(n)}=\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\mathcal{F}^{(0)}=\mu_{n}{\rm e}^{\mathcal{F}^{(0)}(T-n\alpha g_{s})-\mathcal{F}^{(0)}(T)}.$ (4.80)
An important feature of the basic instanton amplitude is the shift by $n$ units of the flat coordinate, as emphasized in Gu:2022sqc , and it corresponds to $n$ units of the RR-flux through the compact $T^{2}$. In addition, it is expressed as a functional of the perturbative free energy. Together with the chain rule of the alien derivative, it allows us to calculate the other instanton amplitudes at multi-instanton orders. For instance, at $2$-instanton order, the other instanton amplitude is
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}\mathcal{F}^{(0)}=\mu_{1}^{2}\left({\rm e}^{\mathcal{F}_{2}^{(0)}-\mathcal{F}^{(0)}}-{\rm e}^{2\mathcal{F}_{1}^{(0)}-2\mathcal{F}^{(0)}}\right).$ (4.81)
At 3-instanton order,
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}\mathcal{F}^{(0)}=$ $\mu_{1}\mu_{2}\left({\rm e}^{\mathcal{F}^{(0)}_{3}-\mathcal{F}^{(0)}}-{\rm e}^{\mathcal{F}_{2}^{(0)}+\mathcal{F}_{1}^{(0)}-2\mathcal{F}^{(0)}}\right),$ (4.82)
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 3}}\mathcal{F}^{(0)}=$ $\mu_{1}^{3}\left({\rm e}^{\mathcal{F}_{3}^{(0)}-\mathcal{F}^{(0)}}-3{\rm e}^{\mathcal{F}_{2}^{(0)}+\mathcal{F}_{1}^{(0)}-2\mathcal{F}^{(0)}}+2{\rm e}^{3\mathcal{F}_{1}^{(0)}-3\mathcal{F}^{(0)}}\right).$ (4.83)
We now check our results explicitly. In terms of alien derivatives and up to 4-instanton order, the Stokes discontinuity is
$\text{disc}_{0}\mathcal{F}^{(0)}=$ $\left(\exp\sum_{n=1}^{\infty}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{n\mathcal{A}}\right)\mathcal{F}^{(0)}-\mathcal{F}^{(0)}$
$=$ $\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}\mathcal{F}^{(0)}+\left(\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}+\frac{1}{2}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}\right)\mathcal{F}^{(0)}$
$+$ $\left(\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{3\mathcal{A}}+\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}+\frac{1}{3!}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 3}}\right)\mathcal{F}^{(0)}$
$+$ $\left(\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{4\mathcal{A}}+\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{3\mathcal{A}}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}+\frac{1}{2}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}+\frac{1}{2}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{2\mathcal{A}}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 2}}+\frac{1}{4!}\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}}^{\raisebox{-3.01389pt}{\scriptsize 4}}\right)\mathcal{F}^{(0)}+\ldots$ (4.84)
We can check (4.84) together with (4.80), (4.81), (4.82), (4.83) numerically. As we have seen in Figs. 4.1 and also from Section 4.2, when $t$ is positive, there is a Stokes ray along the positive real axis due to the Borel singularities $\ell\mathcal{A}_{c}$ with $\ell=1,2,3,\ldots$. We take $t=2\pi/5$, $t=2\pi/7$, calculate numerically the Stokes discontinuity across the positive real axis at $g_{s}=\mathcal{A}_{c}/10$, and subtract successively instanton amplitudes of higher and higher orders. We find that the remainder becomes progressively smaller as shown in Tab. 4.1. We also note that our results only depend on the boundary conditions (4.33) and the HAEs (2.21), and should in fact be valid even when $\theta$ is turned on so that $t$ is complex as well.
$n$ $t=2\pi/5$ $t=2\pi/7$
0 ${\mathsf{i}}\,0.002031$ ${\mathsf{i}}\,0.0003390$
11 $1.5160\times 10^{-8}+{\mathsf{i}}\,4.3236\times 10^{-10} +$$3.4822\times 10^{-8}+{\mathsf{i}}\,1.2783\times 10^{-9}$
22 $8.2065\times 10^{-14}+{\mathsf{i}}\,1.7062\times 10^{-12}$ $3.6852\times 10^{-13}+{\mathsf{i}}\,5.6901\times 10^{-12}$
33 $2.1957\times 10^{-16}-{\mathsf{i}}\,1.4307\times 10^{-17}$ $1.0821\times 10^{-15}-{\mathsf{i}}\,9.6552\times 10^{-17}$
44 $2.4669\times 10^{-21}-{\mathsf{i}}\,3.0214\times 10^{-20} +$$2.4866\times 10^{-20}-{\mathsf{i}}\,2.2110\times 10^{-19}$
55 $4.3291\times 10^{-24}-{\mathsf{i}}\,1.0811\times 10^{-23} +$$4.7115\times 10^{-23}+{\mathsf{i}}\,9.5123\times 10^{-24}$
Table 4.1: Stokes discontinuity of perturbative free energy across the positive real axis after subtracting up to order $n$ instanton amplitudes at $g_{s}=\mathcal{A}_{c}/10$. We use perturbative power series up to 200 terms, and the expected error due to the truncation is of the order $\mathcal{O}(10^{-24})$ in both cases.
We can also work out the Stokes transformation of the partition function. In the $A$-frame, the partition function after the Stokes transformation is
$\mathfrak{S}\mathcal{Z}^{(0)}_{\mathcal{A}}=\exp\mathfrak{S}\mathcal{F}^{(0)}_{\mathcal{A}}=\mathcal{Z}^{(0)}_{\mathcal{A}}\exp\sum_{n=1}^{\infty}\mu_{n}{\rm e}^{-n\mathcal{A}/g_{s}},$ (4.85)
and the reduced partition function is
$\mathcal{Z}_{r,\mathcal{A}}=\mathfrak{S}\mathcal{Z}^{(0)}_{\mathcal{A}}/\mathcal{Z}^{(0)}_{\mathcal{A}}=\exp\sum_{n=1}^{\infty}\mu_{n}{\rm e}^{-n\mathcal{A}/g_{s}}.$ (4.86)
With the Faà di Bruno formula
$\exp\left(\sum_{j=1}^{\infty}x_{j}\frac{t^{j}}{j!}\right)=\sum_{n=0}^{\infty}B_{n}(\{x_{j}\}_{j=1,\ldots,n})\frac{t^{n}}{n!},$ (4.87)
where $B_{n}(\ldots)$ are the Bell polynomials, the reduced partition function in the $A$-frame can be put explicitly in the form (4.56) that we anticipate
$\mathcal{Z}_{r,\mathcal{A}}=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\{j!\mu_{j}\}_{j=1,\ldots,n}\right){\rm e}^{-n\mathcal{A}/g_{s}}.$ (4.88)
Apply again the rule of replacement (4.62), the reduced partition function in a non-$A$-frame reads
$\mathcal{Z}_{r}=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\{j!\mu_{j}\}_{j=1,\ldots,n}\right){\rm e}^{\mathcal{F}(T-n\alpha g_{s})-\mathcal{F}^{(0)}(T)},$ (4.89)
and the non-perturbative partition function after the Stokes transformation is
$\mathfrak{S}\mathcal{Z}^{(0)}=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\{j!\mu_{j}\}_{j=1,\ldots,n}\right)\mathcal{Z}^{(0)}(T-n\alpha g_{s}).$ (4.90)
The result (4.90) also allows us to write down a compact formula for the right hand side of the Stokes discontinuity formula (4.84). With the Faà di Bruno formula for a second time
$\log\left(1+\sum_{j=1}^{\infty}\frac{x_{j}}{j!}t^{j}\right)=\sum_{n=1}^{\infty}\frac{t^{n}}{n!}\sum_{k=1}^{n}(-1)^{k-1}(k-1)!B_{n,k}(\{x_{j}\}_{j=1,\ldots,n-k+1}),$ (4.91)
where $B_{n,k}(\ldots)$ are the incomplete Bell polynomials, we find that
$\text{disc}_{0}\mathcal{F}^{(0)}=\mathfrak{S}\mathcal{F}^{(0)}-\mathcal{F}^{(0)}=\log\mathfrak{S}\mathcal{Z}^{(0)}-\log\mathcal{Z}^{(0)}$ (4.92)
$=$ $\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{k=1}^{n}(-1)^{k-1}(k-1)!{\rm e}^{-k\mathcal{F}^{(0)}}B_{n,k}\left(\{B_{\ell}(\{j!\mu_{j}\}_{j=1,\ldots,\ell}){\rm e}^{\mathcal{F}^{(0)}(T-\ell\alpha g_{s})}\}_{\ell=1,\ldots,n-k+1}\right),$
from which effects of alien derivatives such as (4.81), (4.82), (4.83) and those of even higher orders can be easily read off.
4.5 Non-perturbative proposal
Consider the holomorphic partition function in the large radius frame. When both $t$ and $g_{s}$ are positive, the Borel resummed partition function has a Stokes ray along the positive real axis induced by real instantons with action $\ell t^{2}/2$ ($\ell=1,2,\ldots$), and the Stokes transformation across this Stokes ray is described by (4.90) with $T=t,\alpha=-1$. This tells us how to write down the non-perturbative partition function which includes the contributions of all the real instantons, and which is itself real for positve $t$ and $g_{s}$. It is given by
$\mathcal{Z}^{\text{np}}(t;g_{s};\bm{\sigma})=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}\pm\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n}\right)\mathscr{S}^{(\pm)}\mathcal{Z}(t+ng_{s};g_{s}),$ (4.93)
where $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots)$ is an infinite family of real parameters. Here we can take either of the two lateral Borel resummations $\mathscr{S}^{(\pm)}$, but at the same time we have to choose the same sign $\pm$ inside the Bell polynomials to match it. The two definitions are equivalent due to the Stokes transformation (4.90) as
$\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}+\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n}\right)\mathscr{S}^{(+)}\mathcal{Z}(t+ng_{s};g_{s})$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}+\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n}\right)\mathscr{S}^{(-)}\mathfrak{S}\mathcal{Z}(t+ng_{s};g_{s})$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}\mathscr{S}^{(-)}\mathcal{Z}(t+ng_{s};g_{s})\sum_{m=0}^{n}{n\choose m}B_{m}\left(\left\{j!\,\mu_{j}\right\}_{j=1,\ldots,m}\right)$
$\times B_{n-m}\left(\left\{\sigma_{j}+\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n-m}\right)$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}-\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n}\right)\mathscr{S}^{(-)}\mathcal{Z}(t+ng_{s};g_{s}),$ (4.94)
where $\mu_{j}$ are defined in (4.33) and we have used the law of convolution of Bell polynomials
$\sum_{m=0}^{n}{n\choose m}B_{m}(\{x\})B_{n-m}(\{y\})=B_{n}(\{x+y\}).$ (4.95)
This is also verified numerically in Tabs. 4.2, 4.3, 4.4, 4.5: as higher order instanton corrections are included, both of the two lateral resummations have progressively smaller imaginary parts, and they converge to the same real number. Note that the simplest member of this family is the celebrated medium resummation of the perturbative partition function
$\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})=$ $\mathscr{S}^{\text{med}}\mathcal{Z}(t;g_{s})=\mathscr{S}^{(\pm)}\mathfrak{S}^{\mp 1/2}\mathcal{Z}(t;g_{s})$ (4.96)
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\pm\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,\ldots,n}\right)\mathscr{S}^{(\pm)}\mathcal{Z}(t+ng_{s};g_{s}).$
[…]
$n=0$ $2.95\times 10^{-6}+{\mathsf{i}}\,2.95\times 10^{-8}$ $2.95\times 10^{-6}-{\mathsf{i}}\,2.95\times 10^{-8}$
$n=1$ $2.95\times 10^{-6}-{\mathsf{i}}\,8.26\times 10^{-12}$ $2.95\times 10^{-6}+{\mathsf{i}}\,8.26\times 10^{-12}$
$n=2$ $2.95\times 10^{-6}-{\mathsf{i}}\,5.76\times 10^{-15}$ $2.95\times 10^{-6}+{\mathsf{i}}\,5.76\times 10^{-15}$
$n=3$ $2.95\times 10^{-6}+{\mathsf{i}}\,1.94\times 10^{-19}$ $2.95\times 10^{-6}-{\mathsf{i}}\,1.94\times 10^{-19}$
$n=4$ $2.95\times 10^{-6}+{\mathsf{i}}\,3.26\times 10^{-24}$ $2.95\times 10^{-6}-{\mathsf{i}}\,3.26\times 10^{-24}$
Table 4.2: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})$ at $t=2\pi/5$, $g_{s}=\mathcal{A}_{c}/5$ including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to the truncation is of the order $\mathcal{O}(10^{-24})$.
[…]
$n=0$ $48\times 10^{-4}+{\mathsf{i}}\,1.77\times 10^{-6} 4.4$$48\times 10^{-4}-{\mathsf{i}}\,1.77\times 10^{-6}$
$n=1$ $48\times 10^{-4}-{\mathsf{i}}\,1.08\times 10^{-10} 4.4$$48\times 10^{-4}+{\mathsf{i}}\,1.08\times 10^{-10}$
$n=2$ $48\times 10^{-4}-{\mathsf{i}}\,1.10\times 10^{-15} 4.4$$48\times 10^{-4}+{\mathsf{i}}\,1.10\times 10^{-15}$
$n=3$ $48\times 10^{-4}+{\mathsf{i}}\,2.47\times 10^{-19} 4.4$$48\times 10^{-4}-{\mathsf{i}}\,2.47\times 10^{-19}$
Table 4.3: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})$ at $t=2\pi/3$, $g_{s}=\mathcal{A}_{c}/5$ including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to the truncation is of the order $\mathcal{O}(10^{-20})$.
[…]
$n=0$ $2.95\times 10^{-6}+{\mathsf{i}}\,2.95\times 10^{-8}$ $2.95\times 10^{-6}-{\mathsf{i}}\,2.95\times 10^{-8}$
$n=1$ $2.98\times 10^{-6}+{\mathsf{i}}\,6.20\times 10^{-12}$ $2.98\times 10^{-6}-{\mathsf{i}}\,6.20\times 10^{-12}$
$n=2$ $2.98\times 10^{-6}+{\mathsf{i}}\,1.17\times 10^{-15}$ $2.98\times 10^{-6}-{\mathsf{i}}\,1.17\times 10^{-15}$
$n=3$ $2.98\times 10^{-6}+{\mathsf{i}}\,3.51\times 10^{-19}$ $2.98\times 10^{-6}-{\mathsf{i}}\,3.51\times 10^{-19}$
$n=4$ $2.98\times 10^{-6}+{\mathsf{i}}\,5.11\times 10^{-24}$ $2.98\times 10^{-6}-{\mathsf{i}}\,5.11\times 10^{-24}$
Table 4.4: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s},\bm{\sigma})$ at $t=2\pi/5$, $g_{s}=\mathcal{A}_{c}/5$ with $\bm{\sigma}=(3,2,1,9,\ldots)$ including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to the truncation is of the order $\mathcal{O}(10^{-24})$.
[…]
$n=0$ $48\times 10^{-4}+{\mathsf{i}}\,1.77\times 10^{-6} 4.4$$48\times 10^{-4}-{\mathsf{i}}\,1.77\times 10^{-6}$
$n=1$ $52\times 10^{-4}+{\mathsf{i}}\,1.41\times 10^{-9} 4.4$$48\times 10^{-4}+{\mathsf{i}}\,1.41\times 10^{-9}$
$n=2$ $52\times 10^{-4}+{\mathsf{i}}\,3.21\times 10^{-14} 4.4$$48\times 10^{-4}+{\mathsf{i}}\,2.21\times 10^{-15}$
$n=3$ $52\times 10^{-4}+{\mathsf{i}}\,2.65\times 10^{-19} 4.4$$48\times 10^{-4}-{\mathsf{i}}\,2.65\times 10^{-19}$
Table 4.5: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})$ at $t=2\pi/3$, $g_{s}=\mathcal{A}_{c}/5$ with $\bm{\sigma}=(5,3,7,1,\ldots)$ including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to the truncation is of the order $\mathcal{O}(10^{-20})$.
Let us compare this result with the prosposal of Okuyama and Sakai Okuyama:2018clk summarized in Section 2.3. We can spell out the latter more explicitly. By using the Faà di Bruno formula (4.87), we find
$\phi_{n}(q)=\frac{1}{n!}B_{n}\left(\left\{\frac{{\mathsf{i}}}{2}(j-1)!(-1)^{j}\vartheta_{2}(q^{j})\right\}_{j=1,\ldots,n}\right).$ (4.97)
In the limit of $g_{s}\rightarrow 0$ and $q\rightarrow 1$, by the modular property of Jacobi theta functions
$\vartheta_{2}(q^{\ell})=\vartheta_{2}\left(0;\frac{{\mathsf{i}}\ell g_{s}}{2\pi}\right)=\sqrt{\frac{2\pi}{\ell g_{s}}}\vartheta_{4}\left(0;\frac{2\pi{\mathsf{i}}}{\ell g_{s}}\right)=\sqrt{\frac{2\pi}{\ell g_{s}}}\vartheta_{4}\left({\rm e}^{-4\pi^{2}/(\ell g_{s})}\right),$ (4.98)
(2.36) can be written as
$Z^{\text{OS}}(t;g_{s})=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\vartheta_{4}({\rm e}^{-4\pi^{2}/(jg_{s})})\right\}\right)\mathscr{S}^{(+)}Z^{\text{top}}(t+ng_{s};g_{s}).$ (4.99)
Eq. (4.99) indicates that there are instanton sectors with action $nt^{2}/2+4m\pi^{2}/j$, with $j=1,\ldots,n$ and $m=1,2,\ldots$. They are not observed in the numerical calculations of alien derivatives. This problem can be bypassed by arguing that these additional instanton sectors are in disjoint resurgent structures than that of the perturbative partition function, therefore inaccessible by resurgence analysis of the perturbative sector, just like even and odd instanton sectors in the double-well quantum mechanics model.
More importantly, (4.99) coincides with (4.96) with ++ sign in the leading order with $\vartheta_{4}\approx 1$, but deviates from it by higher order terms. This indicates that (4.99) is not real for positive $t$ and $g_{s}$, i.e. the cancellation of imaginary part is incomplete due to higher order terms in $\vartheta_{4}$. To see this problem more clearly, let us apply the Stokes transformation (4.64),(4.90),
$\mathcal{Z}^{\text{OS}}(t;g_{s})=$ $\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\vartheta_{4}({\rm e}^{-4\pi^{2}/(jg_{s})})\right\}\right)\mathscr{S}^{(-)}\mathfrak{S}Z(t+ng_{s};g_{s})$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}\mathscr{S}^{(-)}Z(t+ng_{s};g_{s})\sum_{m=0}^{n}{n\choose m}B_{m}\left(\left\{-{\mathsf{i}}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}\right)$
$\phantom{===}\times B_{n-m}\left(\left\{\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\vartheta_{4}({\rm e}^{-4\pi^{2}/(jg_{s})})\right\}\right)$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{-\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}(2-\vartheta_{4}({\rm e}^{-4\pi^{2}/(jg_{s})}))\right\}\right)\mathscr{S}^{(-)}\mathcal{Z}(t+ng_{s};g_{s}),$ (4.100)
where we have used again (4.95). Note that the right hand side of (4.100) and (4.99) are not symmetric as in the case of (4.96), which implies that the Borel resummation (4.99) is not real.
[…]
$n=0$ $3.21\times 10^{-5}+{\mathsf{i}}\,4.38\times 10^{-11}$ $3.21\times 10^{-5}-{\mathsf{i}}\,4.38\times 10^{-11}$
$n=1$ $3.21\times 10^{-5}-{\mathsf{i}}\,6.07\times 10^{-21}$ $3.21\times 10^{-5}+{\mathsf{i}}\,3.98\times 10^{-15}$
$n=2$ $3.21\times 10^{-5}+{\mathsf{i}}\,9.40\times 10^{-23}$ $3.21\times 10^{-5}+{\mathsf{i}}\,3.98\times 10^{-15}$
$n=3$ $3.21\times 10^{-5}+{\mathsf{i}}\,9.40\times 10^{-23}$ $3.21\times 10^{-5}+{\mathsf{i}}\,3.98\times 10^{-15}$
Table 4.6: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})$ and the non-perturbative partition function $\mathcal{Z}^{\text{OS}}$ proposed by Okuyama-Sakai, at $t=2\pi$, $g_{s}=\mathcal{A}_{c}/10$, including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to truncation is of the order $\mathcal{O}(10^{-24})$.
[…]
$n=0$ $1.48\times 10^{-7}+{\mathsf{i}}\,9.83\times 10^{-13}$ $1.48\times 10^{-7}+{\mathsf{i}}\,9.84\times 10^{-13}$
$n=1$ $1.48\times 10^{-7}-{\mathsf{i}}\,1.09\times 10^{-20}$ $1.48\times 10^{-7}-{\mathsf{i}}\,1.06\times 10^{-20}$
$n=2$ $1.48\times 10^{-7}-{\mathsf{i}}\,3.17\times 10^{-27}$ $1.48\times 10^{-7}+{\mathsf{i}}\,3.33\times 10^{-22}$
$n=3$ $1.48\times 10^{-7}-{\mathsf{i}}\,3.17\times 10^{-27}$ $1.48\times 10^{-7}+{\mathsf{i}}\,3.33\times 10^{-22}$
Table 4.7: The non-perturbative partition function $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})$ and the non-perturbative partition function $\mathcal{Z}^{\text{OS}}(t;g_{s})$ proposed by Okuyama-Sakai, at $t=4\pi/3$, $g_{s}=\mathcal{A}_{c}/10$, including up to order $n$ instanton corrections. We use perturbative series up to $200$ terms, and the expected error due to truncation is of the order $\mathcal{O}(10^{-27})$.
We verify this numerically in Tabs. 4.6, 4.7: as the order of included instanton corrections increases, the imaginary part of the resummed partition function (4.99) stops decreasing at some point, even though it is still much larger than the expected error due to the truncation of the perturbative power series [^12] This is probably the reason that in (Okuyama:2018clk, , Sec. 5) the two sides of (2.36) differ in numerical tests by a small imaginary part when $N$ is large., in stark contrast to the medium resummed non-perturbative partition function (4.96). [^13] To see the effect of $\vartheta_{4}({\rm e}^{-4\pi^{2}/(jg_{s})})$, one has to choose $t$ appropriately so that the effect of ${\rm e}^{-2\pi^{2}/(jg_{s})}$ is comparable with instanton corrections of the order ${\rm e}^{-t^{2}/(2g_{s})}={\rm e}^{2\pi^{2}\tau^{2}/g_{s}}$.
4.6 Stokes spectrum
In the previous sections, we have focused on the instanton sectors with action $\pm n\mathcal{A}_{c}$ with $n=1,2,3,\ldots$. As showcased in Figs. 4.1, there are additional instanton sectors with different actions which manifest themselves as additional Borel singularities of the perturbative free energy. We will study their properties, especially the Stokes transformation of the perturbative free energy induced by them.
(a) $t=16\pi/3$
(b) $t=4\pi/3\,{\rm e}^{\pi{\mathsf{i}}/6}$
Figure 4.2: Borel singularities of perturbative free energy in the large radius frame. We use perturbative series truncated to 200 terms, and use Padé approximant to mimic the analytic continuation of the Borel transform. The singular points (red) of the approximation would condense to branch cuts if the truncation is pushed to infinity. At $t=16\pi/3$ $(a)$, the branch points (black) have charges $\pm(1,0,0)$, $\pm(1,2,2)$, $\pm(1,-2,2)$, $\pm(1,4,8)$, $\pm(1,-4,8)$. At $t=4\pi/3\,{\rm e}^{\pi{\mathsf{i}}/6}$ $(b)$, the branch points (black) in the 1st and 3rd quadrants have charges $\pm(1,0,0)$, $\pm(1,2,2)$, and the branch points in the 2nd and 4th quadrants have charges $\pm(2,2,1)$.
Additional instanton sectors can be uncovered from the Borel singularities of perturbative free energy in two different ways. We can either increase the value of $t$ or give $t$ a non-trivial $\theta$ angle, as in the left and right panels of Fig. 4.2. In the former case, we find two towers of Borel singularities located at $\mathcal{A}_{\gamma}$ with charge vectors
$\gamma_{(m,\pm,\pm)}=\pm(1,\pm 2m,2m^{2}),\quad m=1,2,\ldots.$ (4.101)
We find that in the large radius frame, the Stokes transformation is
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}_{\gamma_{(m,\pm,\pm)}}}\mathcal{F}^{(0)}=g_{s}^{2}\mathcal{F}^{(1)},$ (4.102)
with $\mathcal{F}^{(1)}$ given explicitly by (4.51) with $\alpha=\pm 1,\beta=\pm m$, slightly modified by the additional factor of $g_{s}^{2}$. In the latter case, we find further Borel singularities, such as the ones with charges
$\gamma_{\pm 2}=\pm(2,2,1).$ (4.103)
In the large radius frame, the Stokes transformation is
$\stackrel{{\scriptstyle\bullet}}{{\Delta}}_{\mathcal{A}_{\gamma_{\pm 2}}}\mathcal{F}^{(0)}=g_{s}^{2}\mu_{2}\exp\left(\mathcal{F}^{(0)}(t+2g_{s})-\mathcal{F}^{(0)}(t)\right),$ (4.104)
which is a bit peculiar, as it is a mixture of the 1-instanton amplitude with $\alpha=2$ and the 2-instanton prefactor $\mu_{2}$. In all these examples, the Stokes constants are identically set to one due to our normalization of instanton amplitudes.
Following the conjecture in Gu:2021ize ; Gu:2022sqc ; Gu:2023mgf , these instanton sectors might correspond to different BPS states – D-brane bound states – in the type IIA string theory compactified on the Calabi-Yau threefold $X_{E}$, with $\gamma$ as the charge vectors. The BPS spectrum of type IIA string in general have wall-crossing phenomenon in the moduli space: across a real codimension one wall of marginal stability characterized by that two integral periods align in the complex plane, the spectrum of BPS states changes. Here the condition of the wall of marginal stability translates to
$\operatorname{Im}\left(\frac{t^{2}/2}{2\pi{\mathsf{i}}t}\right)=\operatorname{Re}t=0\;\Leftrightarrow\;\operatorname{Im}\tau=0.$ (4.105)
In our discussion we have always assumed
$\operatorname{Im}\tau>0\;\Leftrightarrow\;\operatorname{Re}t>0,$ (4.106)
and the wall of marginal stability is the natural boundary of this region.
Figure 4.3: The moduli space and the possible wall of marginal stability of type IIA string compactified on $X_{E}$.
It is possible to cross the wall and consider the region
$\operatorname{Im}\tau<0\;\Leftrightarrow\;\operatorname{Re}t<0,$ (4.107)
as shown in Fig. 4.3. In the holomorphic limit, the free energy $\mathcal{F}_{g}(\tau)$ with $g\geq 2$ are polynomials of $E_{2},E_{4},E_{6}$. We can define $E_{2k}(\tau)$ with $\tau$ in the lower half plane $\mathcal{H}_{-}$ and they are related to the usual $E_{2k}(\tau)$ in the upper half plane $\mathcal{H}_{+}$ by
$E_{2k}(\tau)=E_{2k}(-\tau).$ (4.108)
For instance
$E_{2}(\tau)=1-24\sum_{n=1}^{\infty}\frac{q^{n}}{(1-q^{n})^{2}}=1-24\sum_{n=1}^{\infty}\frac{1/q^{n}}{(1-1/q^{n})^{2}}=E_{2}(-\tau).$ (4.109)
Similarly, we can show
$E_{4}(\tau)=E_{4}(-\tau),\quad E_{6}(\tau)=E_{6}(-\tau).$ (4.110)
In addition, using
$1728\eta(\tau)^{24}=E_{4}(\tau)^{3}-E_{6}(\tau)^{2},\quad\frac{{\rm d}}{2\pi{\mathsf{i}}{\rm d}\tau}\log\eta(\tau)=\frac{E_{2}(\tau)}{24},$ (4.111)
we can also define
$\eta(\tau)=\eta(-\tau),\quad\tau\in\mathcal{H}_{-}.$ (4.112)
This allows us to define the free energy $\mathcal{F}_{g}(\tau)$ with $g\geq 1$ in the regime (4.107). Given the simple identities (4.108), (4.112), we conclude that the instanton sectors in the regime (4.107) are the same as in the regime (4.106), and the wall-crossing for the wall of marginal stability (4.105) is trivial.
5 Conclusion and discussion
In this paper we use the resurgence theory and the previous results on the resurgent structure of topological string theory to study non-perturbative corrections to the topological string model dual to 2d $U(N)$ Yang-Mills on torus, or equivalently the Gromov-Witten theory of an elliptic curve. We find closed form formulas for instanton amplitudes up to arbitrary high instanton orders, based on which we can write down the non-perturbative partition function with contributions from all the real instantons which are in the resurgent structure of the perturbative series. This non-perturbative partition function overcomes limitations of previous proposals and is in particular real for positive modulus and string coupling. We also explore complex instantons and find two infinite towers as well as two additional complex instantons.
Following Gu:2022sqc ; Gu:2023mgf , we expect that all the instanton sectors, including the complex instantons, correspond to BPS D-brane bound states in type II string on $X_{E}$. We have also studied the wall-crossing behavior of these states. It would be interesting to check and verify the spectrum of BPS states in the type II string.
One of the motivations to study the non-perturbative partition function of topological string is to enable precise formulation of the duality between 2d Yang-Mills on torus and topological string for finite $N$, and more generally that of the OSV conjecture. We wish to make progress in this direction in the near future.
In this paper, we only restrict to the study of 2d Yang-Mills with gauge group on torus. For more generic target space of a Riemann surface of arbitrary genus, there have recently been some progress on the nature of the dual string description. It would be interesting to study the non-perturbative aspects of them. Finally, it would also be interesting to study the dual string theory for the 2d Yang-Mills of other gauge groups, for instance of the $B$, $C$, $D$ types in the large $N$ limit.
Acknowledgement
We would like to thank Marcos Marino for participation in the initial stages of the project, and for careful reading of the manuscript. J.G. is supported by the startup funding No. 4007022316 of the Southeast University, and the National Natural Science Foundation of China (General Program) funding No. 12375062. X.W. is supported by the National Natural Science Foundation of China Grants No.12247103.
Index

### v3

@@
 At large but finite $N$, the factorization is only schematic and it needs to be modified by summing over RR fluxes over the finite $T^{2}$ Vafa:2004qa ; Dijkgraaf:2005bp .
-More importantly, while on the one side the partition function of 2d Yang-Mills is non-perturbatively well-defined for any finite $N$, on the other side the topological string partition function is only defined perturbatively.
+More importantly, while on the one side the partition function of 2d Yang-Mills is non-perturbatively well-defined for any finite $N$, on the other side the topological string partition function is only defined perturbatively in the large $N$ expansion.
 One therefore has to find its non-perturbative completion, taking into account appropriate non-perturbative corrections of the order $\mathcal{O}({\rm e}^{-N})$.
+These are spacetime instantons on the string theory side due to D-branes, and they correspond to large $N$ instanton corrections on the 2d YM theory side.
 There have been various attempts to find these non-perturbative corrections Lelli:2002gr ; Matsuo:2004nn ; deMelloKoch:2005rq ; Dhar:2006ru ; Okuyama:2018clk .
 The most promising proposal is given by Okuyama and Sakai Okuyama:2018clk , based on free fermion formulation of the 2d Yang-Mills Minahan:1993np ; Douglas:1993wy .
-However, only the 1-instanton amplitude of the proposal has been tested, and its multi-instanton amplitudes are not only un-tested, but in fact suffer from some limitations.
+However, only the 1-instanton amplitude of the proposal has been tested, and its multi-instanton amplitudes are not only un-tested, but in fact run into various problems of inconsistency, especially for non-vanishing $\theta$ angles.
 In this paper, we make a new proposal for the full non-perturbative partition function for the topological string dual to chiral 2d Yang-Mills, which overcomes the limitations that plague previous proposals.
@@
 In particular, our formulas are valid for non-vanishing $\theta$ as well.
-Based on these results, we propose the non-perturbative partition function,
-$\mathcal{Z}^{\text{np}}(t;g_{s};\bm{\sigma})=\sum_{n=0}^{\infty}\frac{1}{n!}B_{n}\left(\left\{\sigma_{j}\pm\frac{{\mathsf{i}}}{2}j!\sqrt{\frac{2\pi}{g_{s}}}\frac{(-1)^{j}}{j^{3/2}}\right\}_{j=1,2,\ldots,n}\right)\mathscr{S}^{(\pm)}\mathcal{Z}(t+ng_{s};g_{s}),$ (1.1)
-taking into account contributions from all the real instantons, and it is real when both the modulus $t$ and the string coupling $g_{s}$ are positive.
-It involves an infinite family of real parameters $\bm{\sigma}=(\sigma_{1},\sigma_{2},\ldots)$, and in the simplest case $\bm{\sigma}=\bm{0}$, it is just the medium resummation of the perturbative partition function.
-Here $B_{n}(\ldots)$ are complete Bell polynomials, and $\mathscr{S}^{(\pm)}$ denote the lateral Borel resummations.
-Note that one can choose either of the two lateral Borel resummations $\mathscr{S}^{(\pm)}$ (see (4.10) for definition) with the same sign $\pm$ in the Bell polynomials to match it.
-We believe this should be used to formulate the precise duality between 2d Yang-Mills on torus and topological string for finite $N$.
-In addition, we also explore complex instantons, and find two infinite towers of complex instanton sectors.
+Based on these results, we propose the non-perturbative partition function (4.93) which take into account contributions from all real instantons.
+Even though it does not seem to be unique on the surface and depends on an infinite family of real parameters, we further conjecture that there is a canonical choice of these parameters i.e., (4.96), which should be used to formulate the precise duality between 2d Yang-Mills on torus and topological string for large but finite $N$.
+As we mentioned, the non-perturbative corrections that we have consistently included are large $N$ instanton corrections for 2d YM theory, or worldsheet instanton corrections for the dual topological string theory due to D-branes, and these are all real instantons with real instanton actions.
+There could equally be complex instantons with complex instanton actions.
+Even though they do not play an important role in the non-perturbative partition function of the chiral 2d YM theory, they do correspond to important generic D-branes in the dual topological string theory with complex central charges.
+In this paper, we also explore complex instantons, and we find two infinite towers of complex instanton sectors.
 They should correspond to non-trivial BPS D-brane bound states in type II string.
-We also discuss a little the wall-crossing behavior of these instanton sectors.
+We also discuss the wall-crossing behavior of these instanton sectors.
 The structure of the paper is as follows.
@@
 $S=\frac{1}{2g^{2}_{\text{YM}}}\int_{T^{2}}\operatorname{tr}(F\wedge\star F)+\theta\operatorname{tr}F.$ (2.1)
-Here $\operatorname{tr}$ is the trace in the fundamental representation. The partition function can be calculated by summing over representations $R$ of the gauge group Migdal:1975zg ; Rusakov:1990rs
+Here $\operatorname{tr}$ is the trace in the fundamental representation.
+The partition function can be calculated by summing over representations $R$ of the gauge group Migdal:1975zg ; Rusakov:1990rs ; Witten:1991we
 $Z_{N}^{\text{YM}}=\sum_{R}\exp\left(-\frac{1}{2}g_{\text{YM}}^{2}C_{2}(R)+{\mathsf{i}}\,\theta\,C_{1}(R)\right),$ (2.2)
@@
 where
-$e_{\ell}=\frac{B_{2\ell+2}}{2\ell+2}e^{-2\ell}D^{2\ell-1}E_{2\ell+2}(q),$ (2.15)
+$e_{\ell}=\frac{B_{2\ell+2}}{2\ell+2}2^{-2\ell}D^{2\ell-1}E_{2\ell+2}(q),$ (2.15)
 and
@@
 The free energy $\mathcal{F}_{g}(t)$ can then be calculated from the exponential relation (2.12).
-To compute the free energy $\mathcal{F}_{g}(t)$ efficiently, we begin with a high-order, e.g., 2000th order, $q$-series expansion of $e_{\ell}$. Utilizing the recursive relation
+To compute the free energy $\mathcal{F}_{g}(t)$ efficiently, we begin with a high-order, e.g., 1000th order, $q$-series expansion of $e_{\ell}$. Utilizing the recursive relation
 $h_{n}^{\text{top}}=\frac{1}{n}\sum_{l=1}^{n}h_{n-l}^{\text{top}}\frac{l\,e_{l}}{(2l)!},$ (2.19)
 derived from (2.14), the $q$-series expression of $h_{n}^{\text{top}}$ can be computed efficiently.
-Substituting $h_{n}^{\text{top}}$ into the recursive relation (2.13), we are able to obtain the 2000th-order $q$-series expression of $\mathcal{Z}_{n}^{\text{top}}(t)$ for $n\leq 200$.
+Substituting $h_{n}^{\text{top}}$ into the recursive relation (2.13), we are able to obtain the 1000th-order $q$-series expression of $\mathcal{Z}_{n}^{\text{top}}(t)$ for $n\leq 200$.
+[^6] The data and the program to generate these data are put on the website https://github.com/cloud-atlas-47/2dYM-np.
 Since $\mathcal{Z}_{n}^{\text{top}}(t)$ is finitely generated by Eisenstein series, a sufficiently high-order $q$-series expansion allows one, in principle, to recover the modular expression of $\mathcal{Z}_{n}^{\text{top}}(t)$.
+For instance, we obtin the full modular expression for $\mathcal{Z}_{n}^{\text{top}}(t)$ up to $n=60$.
 Finally, the free energy $\mathcal{F}_{g}(t)$ with $g\geq 1$ of topological string theory has the property that it can be uplifted to a non-holomorphic version $F_{g}(t,\bar{t})$, which satisfies the BCOV holomorphic anomaly equations Bershadsky:1993cx ; Bershadsky:1993ta .
@@
 where the holomorphic part $P_{0}(z)$ is the holomorphic free energy $\mathcal{F}_{g}(t)$.
+Using this method, we have calculated the non-holomorphic free energy $F_{g}$ up to $g=60$.
+[^8] The data and the program to generate these data are put on the website https://github.com/cloud-atlas-47/2dYM-np.
 2.3 Non-perturbative proposal
-It was argued in Vafa:2004qa that the factorization (2.4) is only schematic and to make it exact one has to sum over RR-fluxes over the torus, arriving at
+It was pointed out in Vafa:2004qa that the factorization (2.4) was only approximately valid for large but finite $N$, and one had to take into account non-perturbative corrections $\mathcal{O}({\rm e}^{-N})$ by summing over RR-fluxes over the torus, arriving at
 $Z_{N}^{\text{YM}}=\sum_{n\in\mathbb{Z}}\mathcal{Z}_{\text{top}}(t+ng_{s};g_{s})\overline{\mathcal{Z}}_{\text{top}}(t-ng_{s};g_{s}).$ (2.29)
-An even more refined version is proposed in Dijkgraaf:2005bp .
-It is, however, difficult to generalize formulas like (2.29) to cases with large but finite $N$.
-As argued in Okuyama:2018clk , in this case the meaning of (2.29) is unclear: the left hand side is well-defined for finite $N$ and thus finite $g_{s}$, while the right hand side is only defined as asymptotic perturbative power series of $g_{s}$.
-An attempt to solve this problem was made in Okuyama:2019rqn , in which process, a non-perturbative version of topological string free energy was proposed.
+An even more refined version was proposed in Dijkgraaf:2005bp .
+It was later argued in Okuyama:2018clk that even these proposals were flawed: e.g., the left hand side of (2.29) was well-defined for finite $N$ and thus finite $g_{s}$, while the right hand side of (2.29) was only defined as asymptotic perturbative power series of $g_{s}$.
+An attempt to solve this problem was made in Okuyama:2018clk , in which process, a non-perturbative version of topological string free energy was proposed.
 The partition function of the 2d Yang-Mills theory allows a free fermion formulation Minahan:1993np ; Douglas:1993wy
@@
 $=$ $\oint\frac{{\rm d}x}{2\pi{\mathsf{i}}x^{N+1}}\exp\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}x^{k}\begin{cases}\vartheta_{2}({\rm e}^{{\mathsf{i}}k\theta},p^{k}),\quad&\text{even }N,\\ \vartheta_{3}({\rm e}^{{\mathsf{i}}k\theta},p^{k}),\quad&\text{odd }N,\end{cases}$ (2.30)
-where $p={\rm e}^{-g_{s}}$. Let us focus on the simple scenario $\theta=0$ and assume that $N$ is always even. The partition function reduces to
+where $p={\rm e}^{-g_{s}}$. Restricted to the simple scenario of $\theta=0$ and assuming that $N$ is even, the partition function reduces to
 $Z_{N}^{\text{YM}}=\oint\frac{{\rm d}x}{2\pi{\mathsf{i}}x^{N+1}}\prod_{n\in\mathbb{Z}+\frac{1}{2}}\left(1+xp^{\frac{1}{2}n^{2}}\right).$ (2.31)
@@
 $\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})\left(1+\mathcal{O}({\rm e}^{-1/g_{s}})\right).$ (2.35)
-More explicitly, one can show that Okuyama:2018clk
+More explicitly, it was shown in Okuyama:2018clk that
 $\mathcal{Z}^{\text{OS}}(t,g_{s})=\sum_{n=0}^{\infty}\phi_{n}(p)\mathscr{S}^{(+)}\mathcal{Z}^{\text{top}}(t+ng_{s};g_{s}),$ (2.36)
@@
 $\phi_{1}(p)=-\frac{{\mathsf{i}}}{2}\vartheta_{2}(p),$ (2.38)
-and we have
+and one has
 $\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})-\frac{{\mathsf{i}}}{2}\vartheta_{2}(q)\mathcal{Z}^{\text{top}}(t+g_{s};g_{s})+\ldots$ (2.39)
@@
 $\vartheta_{2}(p)=\sqrt{\frac{2\pi}{g_{s}}}\vartheta_{4}({\rm e}^{-4\pi^{2}/g_{s}})\approx\sqrt{\frac{2\pi}{g_{s}}},$ (2.40)
-we further have the approximation,
+one further has the approximation,
 $\mathcal{Z}^{\text{OS}}(t;g_{s})\sim\mathcal{Z}^{\text{top}}(t;g_{s})-\frac{{\mathsf{i}}}{2}\sqrt{\frac{2\pi}{g_{s}}}\mathcal{Z}^{\text{top}}(t+g_{s};g_{s})+\ldots.$ (2.41)
@@
 (2.41) has been verified in detail by resurgence analysis Okuyama:2018clk .
-However, there are several problems with this proposal.
+Let us comment that there are several problems with this proposal.
 First of all, even though this proposal includes corrections from multi-instanton amplitudes to arbitrary high orders, only the 1-instanton amplitude was tested Okuyama:2018clk .
@@
 4.2, 4.3, 4.4, 4.5: as higher order instanton corrections are included, both of the two lateral resummations have progressively smaller imaginary parts, and they converge to the same real number.
-Note that the simplest member of this family is the celebrated medium resummation of the perturbative partition function
+We note that the formula for non-perturbative partition function (4.93) is not unique as it depends on an infinite family of free parameters $\bm{\sigma}$.
+It thus seems that the resurgence analysis does not capure the complete information of non-perturbative corrections.
+We make two comments here.
+Firstly, it is not uncommon that the resurgence analysis fails to uncover the full non-perturbative corrections.
+For instance, in quantum mechanical models whose potentials are symmetric, resurgence analysis does not capture non-perturbative information in the energy spectrum that arise from (quantum mechanically breaking) these symmetries, such as the removal of energy degeneracy in the double-well model with parity symmetry, and the Bloch angle in the cosine model with periodic symmetry (see for instance discussion in Marino2015:instantons ).
+Second, nevertheless, there does not seem to be any additional symmetries in 2d YM which are quantum mechanically broken, parameterized by some order parameter.
+Therefore, we conjecture that we can simply take the simplest member in the family (4.93), which is the celebrated medium resummation of the perturbative partition function
 $\mathcal{Z}^{\text{np}}(t;g_{s};\bm{0})=$ $\mathscr{S}^{\text{med}}\mathcal{Z}(t;g_{s})=\mathscr{S}^{(\pm)}\mathfrak{S}^{\mp 1/2}\mathcal{Z}(t;g_{s})$ (4.96)
@@
 is supported by the National Natural Science Foundation of China Grants No.12247103.
-Index